### v1

# Resource Allocation via Backscatter-Aware Transmit Antenna Selection for Low-PAPR and Ultra-Reliable WSNs

Abstract
This paper addresses the fundamental physical-layer conflict in hybrid Wireless Sensor Networks (WSNs) between high-throughput primary communication and the sensitive power-envelope requirements of passive sensors. We propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework, a novel per-subcarrier selection scheme designed for multi-antenna illuminators operating within a Multi-Dimensional Orthogonal Frequency Division Multiplexing (MD-OFDM) architecture. Unlike conventional signal-to-noise ratio (SNR) centric selection, BC-TAS utilizes a multi-objective cost function to jointly optimize primary link reliability, stabilize the incident energy envelope at passive Surface Acoustic Wave (SAW) sensors, and provide surgical interference nulling toward coexisting victim nodes. To ensure robustness in low-power environments, we integrate a Kalman-based channel smoothing layer that maintains selection stability even under severe Channel State Information (CSI) estimation errors of up to 20%. Numerical evaluations using the non-linear Rapp power amplifier model and IEEE 802.11be dispersive channel models demonstrate that BC-TAS achieves a nearly three-order-of-magnitude reduction in outage probability and a $25\times$ improvement in network energy efficiency compared to standard MU-MIMO baselines. Furthermore, the proposed dual-envelope control mechanism achieves a $2.4$ dB reduction in the transmitter’s Peak-to-Average Power Ratio (PAPR), allowing for a $3$ dB reduction in required input back-off (IBO) while ensuring spectral mask compliance. These results confirm that BC-TAS effectively bridges the “Illuminator’s Gap,” providing active regulation of the RF envelope to ensure hardware linearity and robust sensing-communication coexistence in high-density, power-constrained environments.
I Introduction
The rapid proliferation of Internet of Things (IoT) devices and their integration into large-scale Wireless Sensor Networks (WSNs) has intensified the demand for hybrid communication architectures. Modern industrial and medical environments now require a single active hub—the illuminator—to support high-throughput data transmission to legitimate users while simultaneously energizing and interrogating passive, backscatter-based sensors [Zhang2025_SR, Wu2024_WPCN]. However, this integration creates a fundamental physical-layer conflict: the wideband, high-power waveforms required for reliable primary communication often saturate the sensitive dynamic range of passive tags or cause prohibitive interference to coexisting victim receivers in the same spectral vicinity [Tan2025_Coexist].
Conventional Multiple-Input Multiple-Output (MIMO) techniques, such as Minimum Mean Square Error (MMSE) precoding, are typically optimized for spectral efficiency. However, they are increasingly unsuitable for the next generation of energy-constrained Passive WSNs (PWSNs) due to three critical limitations. First, the requirement for multiple active RF chains imposes a hardware complexity and power consumption floor that is incompatible with green-WSN goals [Zhao2024_Green]. Second, the high Peak-to-Average Power Ratio (PAPR) inherent in multi-stream Orthogonal Frequency Division Multiplexing (OFDM) necessitates significant Power Amplifier (PA) back-off, leading to spectral regrowth and accelerated battery depletion [Park2024_PAPR]. Most importantly, current precoding strategies lack a mechanism to regulate the incident RF envelope perceived by remote passive devices [Wang2024_Sense]. For time-modulated sensors, such as Surface Acoustic Wave (SAW) tags, high-PAPR fluctuations act as a source of “envelope noise,” masking low-power acoustic reflections and degrading sensing fidelity [Liu2025_SAW].
While diversity-oriented selection methods, such as Transmit Antenna Selection (TAS) and Multi-Dimensional OFDM (MD-OFDM), provide a low-complexity alternative with inherent PAPR benefits, they remain largely “sensing-blind” [Garcia2024_IM]. Existing selection literature focuses exclusively on maximizing desired-link Signal-to-Noise Ratio (SNR). Consequently, there is an urgent need for a framework that treats the spatial-frequency resource blocks of an OFDM symbol as a tool for multi-objective field shaping—simultaneously optimizing primary link reliability, stabilizing backscatter energy incidence, and surgically nulling interference toward victim nodes [Ibrahim2025_MultiObj].
To address these gaps, we propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework. This is the first per-subcarrier selection scheme designed to bridge the “Illuminator’s Gap” by actively managing the RF conditions perceived by passive devices. By shifting the selection logic from a purely SNR-centric approach to a multi-objective sensing-aware heuristic, BC-TAS transforms the illuminator to act as a network-wide controller of both signal quality and incident energy stability.
The central innovation of this work is the development of a unified cost function that embeds three interacting channel components—the legitimate link, tag-side incidence, and victim-link interference—into a per-subcarrier decision rule. Furthermore, we leverage the sparsity of MD-OFDM to achieve a dual-regulation effect: minimizing transmitter-side PAPR to ensure PA linearity while simultaneously minimizing the Backscatter Crest Factor (BCF) to ensure tag-side envelope stability. This dual-regulation ensures that the system remains compliant with strict spectral masks, such as those defined in the maturing IEEE 802.11 Ambient Power Communication (AMP) frameworks, even when operating in highly dispersive environments [IEEE80211AMP_Update, IEEE80211bk_2025].
The key contributions of this work are summarized as follows:
  • BC-TAS Framework Design: We introduce a novel multi-objective per-subcarrier selection metric that explicitly balances primary link reliability against the stability of backscatter incidence and the suppression of victim-link interference.
  • Analytical Metric for Sensing Fidelity (BCF): We define and derive the BCF as a novel performance metric to quantify the temporal stability of the RF envelope at the tag. We provide a theoretical bound for BCF reduction under $N_{t}$-order selection diversity, establishing the first formal link between spatial resource allocation and sensing ”envelope noise.
  • Dual-Envelope Regulation: We establish a joint control mechanism that exploits sparse subcarrier activation to simultaneously reduce the transmitter’s PAPR and the tag’s incident BCF, providing the first quantitative link between illuminator antenna selection and backscatter sensing fidelity.
  • Low-Complexity Heuristic: We formulate the BC-TAS decision as a constrained optimization problem and develop a low-complexity greedy solution incorporating recursive Kalman-based channel smoothing to ensure selection stability under imperfect CSI. The algorithm scales linearly with the number of subcarriers, providing a 2.3× execution speedup over traditional MMSE detectors while maintaining robustness to rapid channel fluctuations.
  • Non-linear Hardware Validation: We perform a comprehensive evaluation using the Rapp PA model and IEEE 802.11n/be channel models, demonstrating over two orders of reduction in outage probability and a $25\times$ improvement in system energy efficiency compared to standard SISO and MIMO baselines.
The remainder of this paper is organized as follows. Section II reviews the state-of-the-art in TAS, backscatter sensing, and PAPR-aware transmission. Section III details the system model and the MOFS formulation, including the Kalman-based channel smoothing. Section IV presents the proposed greedy selection algorithm and derives its theoretical diversity and scaling properties. Section V provides extensive simulation results, including BER, BCF, and hardware efficiency metrics. Section VI synthesizes these findings with a theoretical validation of the observed gains and trade-offs. Finally, Section VII concludes the work, followed by the derivation of the selection penalty in Appendix A.
II Related Work
II-A Transmit Antenna Selection (TAS)
TAS remains a cornerstone for low-complexity diversity in modern wireless architectures. Recent research by Kumar et al. (2025) demonstrates that TAS provides significant diversity gains in active RIS-assisted downlink NOMA systems, where the diversity order scales jointly with BS antennas and RIS elements [Kumar2025]. Similarly, Han et al. (2025) emphasize that TAS-based strategies, particularly under statistical Channel State Information (CSI), can effectively improve outage probability and sum throughput without the prohibitive RF-chain costs of full MIMO [Han2025]. However, these classical and 6G-era TAS frameworks primarily optimize internal communication metrics (SNR, Capacity, BER). There remains a critical research gap in applying TAS to manage external physical-layer constraints, such as the RF envelope of a backscatter tag or the interference profile of a sensitive victim node.
II-B OFDM PAPR Reduction and Hardware Linearity
PAPR is the primary driver of PA non-linearity, leading to spectral regrowth and constellation distortion. Contemporary work in Integrated Sensing and Communication (ISAC) has highlighted that PAPR reduction is essential for improving transmitter efficiency and range under non-linear High Power Amplifier (HPA) constraints [Huang2022]. While probabilistic methods like Partial Transmit Sequences (PTS) have been refined to embed side information implicitly to reduce overhead [Goel2022], they still introduce signal processing latency. From a system-level perspective, Komala (2024) demonstrates that inefficient transmission directly translates to reduced network longevity in WSNs [Komala2024]. By mapping these findings to the Rapp model, it becomes evident that an inherently low-PAPR waveform, rather than an algorithmic mitigation layer, is vital for long-lived, energy-efficient Wireless Power and Data Transfer (WPDT).
II-C Index Modulation and MD-OFDM
Index Modulation (IM) and its variants, such as MD-OFDM, have emerged as core 6G enablers for massive connectivity [Li2023]. Unlike classical OFDM, IM exploits the indices of subcarriers or antennas to embed information, providing a sparse activation pattern that can lead to lower PAPR and reduced inter-channel interference [Mgobhozi2023]. Recent advances in permutational IM have even enabled low-complexity, non-coherent detection that is robust under antenna correlation [VilaInsa2025]. However, most IM literature focuses on maximizing spectral efficiency (SE) or reliability (BER). Even when PAPR is addressed, as in multilevel dither signal designs for OFDM-IM [Kim2018], the focus is purely on internal signal integrity. Our work bridges this gap by treating the MD-OFDM index selection as a first-class design variable for shaping the incident energy at a backscatter sensor.
II-D Backscatter Communications and Symbiotic Radio
The evolution of backscatter has moved toward multi-antenna and broadband configurations to improve link reliability in dispersive channels [Zhang2017, Chen2024]. Optimization efforts have traditionally been split: either optimizing tag reflection coefficients for energy-rate trade-offs [Goay2024] or utilizing deep reinforcement learning for beamforming in bistatic setups [Zargari2024]. Recent symbiotic radio research has proposed strategic subcarrier nulling to eliminate direct-link interference in OFDM backscatter [Janjua2025]. Despite these advances, the ”Illuminator’s Gap” persists; few studies investigate using the transmitter’s resource allocation (like BC-TAS) to actively shape the temporal stability of the RF envelope, which is critical for time-modulated sensors like SAW tags.
II-E Coexistence in Passive SAW-based Sensor Networks
As standardization bodies like IEEE 802.11 extend positioning to 320 MHz channels [IEEE80211bk], the need for interference-resilient signaling becomes paramount. Passive SAW sensors represent a unique challenge in this landscape, as they operate without batteries by converting incident RF energy into time-delayed acoustic reflections [Chen2024, Zhang2017]. Unlike active IoT devices, PWSNs are highly susceptible to:
  • RF Envelope Instability: High PAPR signals create power fluctuations that mask the low-power acoustic reflections [Huang2022, Kim2018].
  • Spectral Competition: While frequency-domain nulling has been proposed [Liu2025, Janjua2025], these methods often ignore the hardware non-linearities of the illuminator.
Industry consensus in the IEEE 802.11 AMP group highlights a critical gap: the absence of selection-based mechanisms that utilize the transmitter’s spatial degrees of freedom to optimize the energy envelope at the tag while protecting primary link reliability. The performance of selection-based architectures in industrial settings is inextricably linked to the underlying radio propagation environment. Prior work has introduced Variational Autoencoder (VAE) based models to predict signal-to-interference-plus-noise ratio (SINR) heatmaps within automated Industry 4.0 warehouses, demonstrating how interacting objects within smart infrastructure significantly impact 5G propagation [GuliaVAE2025]. Furthermore, detailed evaluations of 60 GHz millimeter-wave connectivity in these environments reveal that non-line-of-sight (nLOS) paths created by metallic shelving and autonomous handling agents introduce complex multipath fading effects [Gulia60GHz2023]. To manage these dynamics without high computational costs, stabilized symbolic regression has been utilized to derive compact, interpretable expressions for block error rate (BLER) prediction, offering neural-network-level accuracy suitable for real-time embedded deployment [GuliaSymbolic2024]. These studies underscore the necessity for low-complexity, environment-aware selection mechanisms like BC-TAS that can adapt to the complex spatial profiles of automated warehouses.
Recognizing these coexistence challenges, the industrial shift toward sustainable IoT is evidenced by the formation of the IEEE 802.11 AMP Task Group (TGbp). According to the project authorization scope, the group aims to enable battery-free station (STA) operation through energy harvesting and ultra-low-complexity signaling [IEEE80211AMP_PAR]. A primary challenge identified in the group’s functional requirements is the coexistence of these passive-like devices with legacy high-throughput STAs [IEEE80211AMP_Update]. While the task group explores physical-layer modifications for 2.4 GHz and sub-1 GHz bands, current proposals primarily focus on frequency-domain resource allocation. Our work, BC-TAS, complements these efforts by introducing a spatial-domain selection mechanism that addresses the hardware non-linearities and peak-envelope requirements that are not yet fully specified in the maturing AMP framework.
III System Model and BC-TAS Formulation
We consider a multi-antenna WPDT network operating in a dispersive indoor environment. The network consists of a sensor hub (Tx) equipped with $N_{t}$ antennas, an active legitimate receiver (Rx) with $N_{r}$ antennas, a passive SAW sensor (Tag), and a sensitive victim node (Victim).
III-A MD-OFDM and Incident Signal Model
The Tx utilizes a MD-OFDM waveform with $N_{c}$ subcarriers. To maintain high hardware efficiency and linearity, a TAS architecture is adopted where a single antenna $j_{k}\in\{1,\dots,N_{t}\}$ is activated for subcarrier $k$. Let $\mathbf{J}=[j_{1},j_{2},\dots,j_{N_{c}}]$ denote the antenna selection map. The incident signal at the passive Tag, $y_{T}(t)$, is the sum of the selected antenna contributions:
$y_{T}(t)=\frac{1}{\sqrt{N_{c}}}\sum_{k=1}^{N_{c}}\mathbf{h}_{j_{k},k}^{(T)}X_{k}e^{j2\pi k\Delta ft}+n_{T}(t)$ (1)
where $\mathbf{h}_{j_{k},k}^{(T)}$ is the channel coefficient between the selected antenna for subcarrier $k$ and the Tag. In the SAW-WSN context, the Tag’s sensing fidelity depends on the temporal consistency of the RF envelope to prevent echo masking. This stability is quantified via the BCF, defined as a functional of the selection map $\mathbf{J}$:
$\text{BCF}(\mathbf{J})\triangleq\frac{\max_{t}|y_{T}(t)|^{2}}{\mathbb{E}[|y_{T}(t)|^{2}]}=\frac{\max_{t}\left|\sum_{k=1}^{N_{c}}\mathbf{h}_{j_{k},k}^{(T)}X_{k}e^{j2\pi k\Delta ft}\right|^{2}}{\sum_{k=1}^{N_{c}}\mathbb{E}[|\mathbf{h}_{j_{k},k}^{(T)}X_{k}|^{2}]}$ (2)
A BCF of unity represents a perfectly flat incident envelope, which maximizes the signal-to-clutter ratio for SAW reflections.
III-B PA Non-linearity and Spectral Regrowth
The transmitter chain is constrained by the non-linear behavior of the PA. We employ the Rapp model to characterize AM/AM distortion, where the output amplitude $A_{\mathrm{out}}(t)$ relates to the input $A_{\mathrm{in}}(t)$ as:
$A_{\mathrm{out}}(t)=\frac{GA_{\mathrm{in}}(t)}{\left(1+\left(\frac{GA_{\mathrm{in}}(t)}{A_{\mathrm{sat}}}\right)^{2p}\right)^{\frac{1}{2p}}}$ (3)
where $G$ is the small-signal gain and $p$ is the smoothness factor. Higher BCF at the tag and PAPR at the transmitter necessitate larger Input Back-Off (IBO), reducing Energy Efficiency (EE) and increasing spectral leakage into the Victim’s band.
III-C Kalman-Based Channel Smoothing
While the signal characterization in Section III-A assumes known coefficients, practical backscatter links are plagued by estimation errors. To enhance the robustness of the antenna selection logic against measurement noise and estimation errors in $\mathbf{H}_{L}$, we implement a recursive scalar Kalman Filter. Given that adjacent subcarriers in MD-OFDM exhibit high frequency correlation, the filter tracks the latent power gain state of each antenna link across the subcarrier index $k$.
For each transmit antenna $j\in\{1,\dots,N_{t}\}$, the state equation and the measurement equation are defined as:
$x_{j,k}=x_{j,k-1}+w_{k},\quad w_{k}\sim\mathcal{N}(0,Q)$ (4)
$z_{j,k}=x_{j,k}+v_{k},\quad v_{k}\sim\mathcal{N}(0,R)$ (5)
where $x_{j,k}=|H_{L,j,k}|^{2}$ represents the true power gain, $z_{j,k}$ is the noisy measurement derived from the estimated CSI, and $Q$ and $R$ are the process and measurement noise covariances, respectively.
The recursive update for the smoothed gain estimate $\hat{x}_{j,k}$ and the error covariance $P_{j,k}$ is performed as follows:
1) Prediction Stage:
$\hat{x}_{j,k}^{-}=\hat{x}_{j,k-1}$ (6)
$P_{j,k}^{-}=P_{j,k-1}+Q$ (7)
2) Correction Stage:
$K_{k}=\frac{P_{j,k}^{-}}{P_{j,k}^{-}+R}$ (8)
$\hat{x}_{j,k}=\hat{x}_{j,k}^{-}+K_{k}(z_{j,k}-\hat{x}_{j,k}^{-})$ (9)
$P_{j,k}=(1-K_{k})P_{j,k}^{-}$ (10)
The resulting smoothed gain $G_{L,\text{smooth}}=\hat{x}_{j,k}$ is subsequently utilized in the multi-objective cost function (Eq. 11), ensuring that the selection process prioritizes stable, high-gain paths over transient noise spikes.
III-D Multi-Objective Field Shaping (MOFS) Framework
The BC-TAS framework treats antenna selection as a real-time field-shaping problem. For each subcarrier $k$, the optimal antenna index $j_{k}^{\star}$ is determined by minimizing the MOFS cost function. To ensure stability against measurement noise, the algorithm utilizes the Kalman-smoothed gains $\hat{x}_{j,k}$ derived in Section III-C:
$j_{k}^{\star}=\min_{j\in\{1,\dots,N_{t}\}}\left[\frac{\mathcal{P}_{tx}}{\hat{x}_{j,k}+\epsilon}+\lambda_{T}\|h_{j,k}^{(T)}-\bar{H}_{T}\|^{2}+\lambda_{V}\|h_{j,k}^{(V)}\|^{2}\right]$ (11)
where $\lambda_{T}$ and $\lambda_{V}$ are regularization weights. The first term prioritizes primary link SNR. The second term performs field flattening by penalizing deviations from the average incident gain $\bar{H}_{T}$, effectively suppressing the frequency-selective spikes that drive high BCF. The third term provides spatial nulling to protect the Victim node.
III-E Sensing-Centric Evaluation Metrics
To bridge the gap between communication and sensing, we define metrics that quantify the ”visibility” of the passive tag:
  • Sensor Dynamic Range (SDR): The ratio of the detectable SAW reflection power $P_{\text{SAW}}$ to the effective background floor, including thermal noise $\sigma_{n}^{2}$ and the residual envelope fluctuations $P_{B}$ managed by $\lambda_{T}$:
    $\text{SDR}_{\mathrm{dB}}=10\log_{10}\left(\frac{P_{\text{SAW}}}{\sigma_{n}^{2}+\mathbb{E}[P_{B}|\lambda_{T}]}\right)$ (12)
  • Harvesting Efficiency ($\eta_{H}$): The probability that the incident power on subcarrier $k$, $P_{in}(k)$, satisfies the SAW activation threshold $P_{\text{th}}$.
IV Greedy Selection and Theoretical Properties
Lemma 1.
(Diversity Order and Scaling): Under i.i.d. Rayleigh fading, the BC-TAS policy preserves a diversity order of $L=N_{t}$ for the primary link while ensuring the expected interference power at the Tag, $\mathbb{E}[P_{T}]$, scales as $\mathcal{O}(1/N_{t})$.
Proof.
Let $Z=\min_{j}\mathcal{L}(j,k)$. Since the channels are independent, the selection logic effectively samples the lower tail of the interference distribution. As $N_{t}\to\infty$, the probability of finding an antenna that resides in the spatial null of both the Tag and Victim link increases exponentially. The expected interference at the Tag is the mean of the minimum of $N_{t}$ exponential variables, which is $\sigma_{T}^{2}/N_{t}$. A rigorous derivation of this scaling law and the associated penalty function $\Phi(\lambda_{T})$ is provided in Appendix A. ∎
While BC-TAS optimizes the sensing environment, prioritizing the Tag and Victim links introduces a selection penalty $\Delta$ on the primary communication link. This penalty represents the ratio of the achieved legitimate-link gain to the maximum possible TAS gain. In the limit of high interference suppression ($\lambda_{T},\lambda_{V}\gg 1$), the primary link’s diversity gain transitions toward a sample mean. The closed-form derivation of $\Delta$ and its asymptotic behavior as a function of $N_{t}$ are detailed in Appendix A.
While an exhaustive search for the optimal antenna-to-subcarrier mapping $j_{k}^{\star}$ ensures the global minimum of the cost function $\mathcal{L}_{\text{MOFS}}$, its complexity grows exponentially with the number of subcarriers $N_{c}$. To enable real-time implementation on hardware with limited processing power, we propose a low-complexity greedy heuristic based on MOFS.
IV-A Algorithmic Formulation
The BC-TAS algorithm decomposes the multi-objective optimization problem into parallel per-subcarrier decisions. The key innovation is the inclusion of a target gain reference, $\bar{H}_{T}$, which guides the selection toward a flat frequency response at the Tag’s coordinate.
Algorithm 1 BC-TAS Greedy Selection with Field Flattening
0: Channels $\{\mathbf{h}_{j,k}^{(L)},\mathbf{h}_{j,k}^{(T)},\mathbf{h}_{j,k}^{(V)}\}$, Weights $\{\lambda_{T},\lambda_{V}\}$, $N_{c},N_{t}$
0: Optimal Antenna Map $\mathbf{J}=\{j_{1}^{\star},\dots,j_{N_{c}}^{\star}\}$ Initialization : Set $\epsilon=10^{-6}$
1:// Step 1: Compute Target Gain for Field Flattening
[…]
3:for $k=1$ to $N_{c}$ do
4:  // Step 2: Evaluate MOFS Cost for all candidate antennas
5:  for $j=1$ to $N_{t}$ do
6:   $\mathcal{L}_{\text{MOFS}}(j,k)=\frac{1}{\|\mathbf{h}_{j,k}^{(L)}\|^{2}+\epsilon}+\lambda_{T}\|\mathbf{h}_{j,k}^{(T)}-\bar{H}_{T}\|^{2}+\lambda_{V}\|\mathbf{h}_{j,k}^{(V)}\|^{2}$
7:  end for
8:  // Step 3: Select antenna minimizing the composite field deviation
9:  $j_{k}^{\star}=\operatorname*{arg\,min}_{j}\,\mathcal{L}_{\text{MOFS}}(j,k)$
10:end for
11:return $\mathbf{J}$
IV-B Complexity Analysis and Hardware Overhead
The computational efficiency of the BC-TAS framework is a critical factor for its deployment in high-speed industrial WSNs. The complexity is analyzed as follows:
  • Pre-computation Overhead: The calculation of the target gain $\bar{H}_{T}$ requires a double summation over $N_{c}$ and $N_{t}$. This introduces a one-time overhead of $\mathcal{O}(N_{c}N_{t})$ additions per frame. Since this value is constant for all subcarriers within a single coherence interval, it is computed outside the main selection loop, adding negligible latency compared to the symbol duration.
  • Time Complexity: The main selection process consists of a nested loop structure. For each of the $N_{c}$ subcarriers, the algorithm performs $N_{t}$ scalar cost evaluations. Each evaluation involves only basic arithmetic operations (addition, multiplication, and a single reciprocal). Consequently, the total time complexity is $\mathcal{O}(N_{c}N_{t})$. This is significantly lower than standard MIMO-MMSE precoding, which typically scales with $\mathcal{O}(N_{c}N_{r}^{3})$ due to the requirement for matrix inversions at every resource block.
  • Space Complexity: The algorithm is highly memory-efficient, requiring only $\mathcal{O}(N_{t})$ storage to hold the instantaneous channel power values for the current subcarrier evaluation. The storage of $\bar{H}_{T}$ requires only a single scalar register, making BC-TAS ideal for implementation on low-power FPGA or ARM-based IoT gateways.
Remark: In practical hardware, the $\mathcal{O}(N_{c}N_{t})$ operations can be further accelerated through parallelization, as the decision for subcarrier $k$ is independent of subcarrier $k+1$, allowing for a fully pipelined architecture.
IV-C Optimality Gap and Heuristic Justification
The greedy approach is locally optimal for each subcarrier. While it does not explicitly perform inter-subcarrier power loading (like Water-filling), the inherent sparsity of the MD-OFDM framework ensures that the resulting PAPR is statistically bounded. As shown in Section V, the performance gap between the greedy BC-TAS and the exhaustive search remains below $0.5$ dB in the high-SNR regime, justifying the use of the low-complexity heuristic for practical deployments.
V Simulation and Results
This section details the simulation environment, defines the key parameters used, and presents the performance metrics of the proposed Backscattered Constrained Transmit Antenna Selection Multi-Dimensional Orthogonal Frequency Division Multiplexing (BC-TAS MD-OFDM) model compared to baseline schemes.
V-A Simulation Setup
The system performance was evaluated using Monte Carlo simulations. The core parameters are summarized in Table I. To ensure relevance to emerging standards, the channel environment is modeled based on the IEEE 802.11be (Wi-Fi 7) TGbe indoor specifications, utilizing Multipath Fading Models (TGn A–F) to represent varying degrees of frequency selectivity in warehouse environments.
To characterize hardware-induced distortions, we incorporate a non-linear Rapp PA model at the transmitter. The AM/AM conversion follows the smoothness factor $p=2$, which represents a typical Solid State Power Amplifier (SSPA). This model is critical for evaluating the impact of high-PAPR and high-BCF waveforms on spectral regrowth and Legitimate-link Bit Error Rate (BER).
Simulations were performed on a high-performance computing node equipped with an NVIDIA Tesla V100-PCIE-16GB GPU and Intel Xeon processors, utilizing the NVIDIA CUDA 12.3 compiler driver. Execution times reported in Table II represent the total duration required to process $10^{7}$ bits across an SNR range of 0 to 28 dB.
TABLE I: Key Simulation Parameters
Parameter Category Parameter Value
System and Channel Model
System Scheme Proposed BC-TAS MD-OFDM
Baseline Schemes Conventional TAS / MMSE $N_{T}=4,N_{R}=4$
Standard Context IEEE 802.11be EHT Physical Layer
Channel Model Indoor Multipath Fading TGn Models A–F
Fading Distribution Rayleigh Independent Fading Paths
CSI Assumption P-CSI / I-CSI Perfect / Estimation Error $\sigma_{e}^{2}$
OFDM and PHY Layer
Subcarrier Count ($N$) $N_{\mathrm{SC}}$ 256 / 1024
Cyclic Prefix Length ($N_{\mathrm{CP}}$) $L_{\mathrm{CP}}$ 32 samples
Modulation Scheme $M$-QAM 16-QAM / 64-QAM
Total Simulated Bits $N_{\mathrm{bits}}$ $10^{7}$ bits
Non-Linearity and Field Shaping
PA Model Rapp Model $p=2$ (Smoothness factor)
MOFS Weights $\lambda_{T},\lambda_{V}$ $[0.1,0.5,1.0]$
Tag/Victim Distances $d_{L},d_{T},d_{V}$ 10 m, 2 m, 5 m
Kalman Process Noise ($Q$) $Q_{KF}$ $1\times 10^{-4}$
Kalman Meas. Noise ($R$) $R_{KF}$ $1\times 10^{-2}$
Monte Carlo Trials $N_{trials}$ 1,000
V-B Comparative Analysis with State-of-the-Art (SOTA) Baselines
To validate the competitive standing of the proposed BC-TAS framework, we benchmarked its performance against four distinct SOTA categories: fixed-diversity baselines [SOTA_2_MU_STBC], low-complexity selection methods [SOTA_1_MMSE], theoretical upper bounds (OES) [SOTA_3_Exhaustive_Search], and backscatter-specific multi-antenna contexts [SOTA_4_Backscatter_Multi_Antenna].
Figure 1: BER comparison of the proposed BC-TAS framework against SOTA baselines in TGn Model F.
As shown in the simulation results in Fig. 1, the proposed BC-TAS significantly outperforms the industry-standard Norm-Based Antenna Selection (NBAS) [SOTA_5_NBAS]. At an average SNR of 10 dB, BC-TAS achieves a BER of $\approx 3.2\times 10^{-4}$, whereas NBAS remains at $\approx 2.6\times 10^{-2}$. This order-of-magnitude improvement is attributed to our model’s ability to exploit frequency-selective diversity rather than relying on a static norm-based metric. The proposed scheme achieves a near-optimal slope while maintaining robustness to high CSI error ($20\%$).
V-C Impact of CSI Imperfection and Robustness Analysis
A critical requirement for practical backscatter deployments is resilience to CSI estimation errors, which are prevalent due to the low-power nature of the tag. We evaluate the robustness of the proposed BC-TAS framework under three regimes: Perfect CSI (P-CSI), 5% estimation error (csi_error_low=0.05csi\_error\_low=0.05), and a severe 20% error condition (csi_error_high=0.2csi\_error\_high=0.2).
As illustrated in Fig. 1, the BC-TAS (Perfect) configuration defines the heuristic’s lower bound, tracking closely with the MD-OFDM (Max Gain) ceiling until approximately 10 dB SNR. In this regime, the system effectively exploits the full spatial-frequency diversity available in the TGn Model F environment.
  • Moderate Imperfection (5% Error): Under a 5% CSI estimation error, the BC-TAS model exhibits a negligible SNR penalty of approximately 1.5 dB at a target BER of $10^{-3}$. Notably, the diversity slope remains steep, indicating that the multi-objective selection metric is still capable of identifying favorable subcarrier-antenna mappings despite minor phase and magnitude perturbations.
  • Severe Imperfection (20% Error): At 20% error, the system encounters a performance floor at a BER of $\approx 1.8\times 10^{-2}$. While the high error rate significantly degrades the selection accuracy, the proposed BC-TAS (20% Error) remarkably maintains a lower BER than the Norm-Based TAS and MMSE 4x4 baselines in the low-to-mid SNR regions (0–8 dB).
The distinct robustness of the proposed model over NBAS is attributed to the fundamental difference in their selection logic. While the performance of NBAS [SOTA_5_NBAS] collapses under severe 20% CSI error—often erroneously selecting a ”weak” antenna path due to noisy norm estimations—the multi-objective BC-TAS remains more resilient. By considering the average channel state across all 256 subcarriers through its heuristic metric, BC-TAS provides a ”smoothing” effect against noise. This prevents the selection of consistently poor paths, ensuring that the probability of selecting a deep-fade antenna across the entire bandwidth is significantly minimized, thereby maintaining a reliable communication link even under high CSI aging or estimation error.
V-D Computational Efficiency vs. Performance Ceiling
A critical finding of this study is the relationship between the proposed heuristic and the Optimal Exhaustive Search (OES) baseline. While OES [SOTA_3_Exhaustive_Search] defines the absolute performance ceiling—achieving zero BER at 8 dB—it requires a per-carrier exhaustive search across all $N_{sc}$ subcarriers.
[…]
TABLE II: Computational Complexity and Execution Time Benchmark
Scheme Complexity Order Selection Logic Exec. Time (s)
MMSE 4x4 [SOTA_1_MMSE] $\mathcal{O}(N_{sc}N_{R}^{3})$ Spatial Multiplexing 181.69
NBAS [SOTA_5_NBAS] $\mathcal{O}(N_{T}N_{R})$ Max. Norm (Fixed) 166.15
STBC 4x1 [SOTA_2_MU_STBC] $\mathcal{O}(N_{sc}N_{R})$ Orthogonal Coding 102.87
MD-OFDM (OES) [SOTA_3_Exhaustive_Search] $\mathcal{O}(N_{sc}N_{T}N_{R})$ Exhaustive Search 77.37
Proposed BC-TAS $\mathbf{\mathcal{O}(N_{T}N_{R})}$ Heuristic Multi-Objective 79.16
V-E BER Performance and Environmental Robustness
The primary performance metric is the BER versus the SNR. The diversity gain provided by the BC-TAS algorithm is demonstrated by the steeper slope of the BER curve compared to the MMSE baseline.
Figure 2: BER performance of the proposed BC-TAS MD-OFDM across various IEEE 802.11TGn indoor environments (Models A–F).
To validate the model’s robustness, performance is analyzed across the full suite of TGn channel models (A–F), as illustrated in Fig. 2. We observe that while the system maintains high reliability in all scenarios, the environment significantly impacts the diversity order. Model A (Flat/Small Office) exhibits the best performance due to the lack of severe frequency selectivity.
In contrast, as we move toward Model F (Large Open Space) with higher delay spreads, the increased frequency-selective fading introduces more frequent deep fades. However, the MD-OFDM selection diversity effectively mitigates these effects, achieving a $10^{-3}$ BER at approximately 12–14 dB SNR across all dispersive models. This demonstrates the framework’s ability to maintain a consistent sensing-communication link regardless of the physical indoor layout.
V-F Outage Probability and Diversity Gain Analysis
A critical bottleneck for WSNs is the prevalence of deep fading in indoor environments, which leads to link outages and energy-inefficient retransmissions. To evaluate the robustness of BC-TAS, we analyze the Cumulative Distribution Function (CDF) of the instantaneous SINR, $\gamma$, in a frequency-selective TGn Model D environment.
The outage probability is defined as $P_{\mathrm{out}}=P(\gamma<\gamma_{\mathrm{th}})$, where the target threshold is set to $\gamma_{\mathrm{th}}=7$ dB. This threshold represents the minimum operating SINR required for reliable QPSK demodulation and successful packet decoding in standard low-power protocols.
Figure 3: Outage Probability CDF in TGn Model D at an average $\mathrm{SNR}=16$ dB.
The simulation results, illustrated in Fig. 3 on a semi-logarithmic scale, reveal that the proposed BC-TAS scheme significantly reshapes the SINR distribution. While conventional random selection closely tracks the Theoretical Rayleigh (SISO) baseline—exhibiting an outage probability of approximately 0.10580.1058—the BC-TAS scheme suppresses the lower tail of the distribution through high-order selection diversity. The proposed BC-TAS scheme demonstrates a significantly steeper diversity slope compared to the Theoretical Rayleigh (SISO) and conventional random selection baselines. Specifically, at the $7$ dB threshold, BC-TAS achieves an outage probability of $0.000175$, representing a $604.77\times$ reduction in link outages.
TABLE III: Diversity Gain Analysis ($\mathrm{SNR}=16$ dB)
Antennas ($N_{T}$) Outage Prob. ($P_{\mathrm{out}}$) Rel. Gain
1 (SISO) $1.05\times 10^{-1}$ Baseline
2 $\approx 1.12\times 10^{-2}$ $9.4\times$
3 $\approx 1.25\times 10^{-3}$ $84.6\times$
4 (BC-TAS) $\mathbf{1.75\times 10^{-4}}$ $\mathbf{604.7\times}$
As detailed in Table III, this massive improvement stems from the transition from a diversity order of $L=1$ (SISO) to an achieved diversity order of $L=N_{T}=4$. The table demonstrates that the reliability gain grows exponentially as additional spatial paths are exploited. In Rayleigh fading, the outage probability for selection diversity scales as $(P_{\mathrm{out,SISO}})^{N_{T}}$; thus, the observed three-order-of-magnitude reduction is consistent with theoretical expectations for a four-antenna selection system. By effectively ”ignoring” subcarriers in deep fades, BC-TAS ensures that the signal remains above the decoding threshold almost $99.99\%$ of the time, providing a robust backscatter-aware link without increasing hardware complexity.
V-G Backscatter Crest Factor (BCF) and Link Stability
While BER and outage probability characterize the primary link, the stability of the reflected field at the passive tag is governed by the BCF. As derived in Lemma 1, the BC-TAS framework utilizes a field-flattening cost term to minimize the temporal peaks of the backscatter signal.
To ensure statistical rigor, we performed $1,000$ Monte Carlo realizations for each antenna configuration ($N_{t}\in\{2,4,8,16,32\}$). Furthermore, to mitigate the impact of imperfect CSI and prevent erratic antenna switching, we integrated a scalar Kalman Filter to smooth the estimated channel gains before selection.
Figure 4: Average BCF vs. $N_{t}$ across $1,000$ Monte Carlo trials in TGn Model B.
As illustrated in Fig. 4, the proposed BC-TAS (Kalman) model exhibits a superior downward trend in BCF compared to all baselines. At $N_{t}=32$, BC-TAS achieves a BCF of $\approx 14.0$ dB, representing a $1.1$ dB improvement over the OES and a $2.5$ dB improvement over the industry-standard NBAS. The proposed BC-TAS (Kalman) achieves a monotonic reduction in signal peaks, outperforming OES and SOTA baselines.
This gap is significant because OES [SOTA_3_Exhaustive_Search], while optimal for the primary receiver’s BER, is ”tag-blind” and selects antennas that create constructive interference peaks at the tag’s location. In contrast, BC-TAS intelligently sacrifices marginal SNR gains to maintain a ”flatter” reflected field. The inclusion of the Kalman filter is vital here; by providing a smoothed estimate of the channel state, it prevents the high BCF values associated with rapid, noise-driven antenna switching across adjacent subcarriers.
V-H PA Distortion Analysis
To evaluate the impact of non-linear hardware impairments, we employ the Rapp PA model with a smoothness factor $p=2$ and a saturation voltage $V_{sat}=1$. The distortion is quantified through Error Vector Magnitude (EVM) and spectral compliance with the IEEE 802.11n mask.
Figure 5: Smoothed EVM performance vs. IBO highlighting the 16-QAM compliance threshold.
As illustrated in the smoothed analysis in Fig. 5, the proposed BC-TAS maintains a consistent linearity advantage over the MMSE MU-MIMO baseline across the entire operational range. While the MMSE baseline remains significantly above the 16-QAM EVM limit of $5.6\%$ throughout the saturation-to-linear transition, the proposed BC-TAS exhibits a steep descent toward compliance as the IBO increases. Specifically, BC-TAS approaches the required $5.6\%$ threshold at an IBO of approximately $13$ dB, demonstrating a significant reduction in constellation distortion compared to MU-MIMO.
Furthermore, the indoor spectral compliance analysis in Fig. 8 confirms that the BC-TAS architecture successfully suppresses out-of-band emissions. Even at an IBO of $10$ dB, the normalized PSD remains strictly within the 802.11n spectral mask limits, with the selection diversity effectively mitigating the spectral regrowth typically associated with high-PAPR OFDM signals. By operating the PA with less back-off while maintaining signal integrity, BC-TAS provides a double benefit: it maximizes the average transmit power available for the backscatter tag’s energy harvesting while simultaneously ensuring robust coexistence with legacy primary receivers in the 2.4/5 GHz bands.
V-I Hardware Performance and Complexity Analysis
The hardware feasibility of the proposed BC-TAS is evaluated by analyzing its robustness against PA non-linearities and its overall energy footprint compared to conventional MU-MIMO baselines.
V-I1 PAPR Reduction and Statistical Analysis
The PAPR is a critical bottleneck for PA efficiency in multi-antenna systems. We evaluate the statistical distribution of these peak power events using the Complementary Cumulative Distribution Function (CCDF), which defines the probability that the PAPR of an OFDM symbol exceeds a specific threshold $\gamma_{0}$.
Figure 6: PAPR Reduction Analysis: Proposed BC-TAS vs. Conventional MMSE MU-MIMO highlighting the selection diversity gain.
As illustrated in Fig. 6, the proposed BC-TAS architecture exhibits a profound statistical advantage over conventional multiplexing-based methods. While both systems converge at low PAPR thresholds where average signal behavior dominates, they diverge significantly in the high-power tail region. At a clipping probability of $10^{-2}$ (1%), the proposed BC-TAS maintains a PAPR of approximately 7.1 dB, whereas the conventional MMSE MU-MIMO baseline reaches 9.5 dB.
This significant 2.4 dB reduction in peak power is a direct consequence of the selection-based architecture. By leveraging the backscatter link to intelligently select subcarrier-antenna mappings that minimize constructive interference, BC-TAS effectively ”truncates” the probability of extreme peak events. This reduction in the PAPR tail is the primary driver for the improved hardware linearity and energy harvesting efficiency observed in subsequent analyses.
V-I2 Energy Efficiency (EE)
Energy Efficiency is defined as the ratio of spectral efficiency to the total power consumption, including RF chain activation and computational costs ($P_{RF}+P_{SEL}$).
Figure 7: System Energy Efficiency vs. PA Operating Point (Input Back-Off).
TABLE IV: Hardware Efficiency and Complexity Comparison ($N_{t}=4$)
Metric Proposed BC-TAS MMSE MU-MIMO Gain
PAPR (@ $10^{-3}$) 10.1 dB 11.2 dB 1.1 dB
[…]
RF Chain Usage 1 4 75% Save
Computational Cost $\mathcal{O}(N_{t})$ $\mathcal{O}(N_{t}^{3})$ Significant
Our results in Fig. 7 demonstrate that BC-TAS provides a massive energy efficiency advantage, peaking at 0.075 bits/Hz/mW. In contrast, the MMSE baseline remains below 0.01 bits/Hz/mW due to the simultaneous activation of $N_{t}$ RF chains and high-complexity matrix inversions.
V-I3 Spectral Mask Compliance
Using the Rapp non-linearity model ($\rho=2$), we evaluate spectral regrowth against the IEEE 802.11n 20 MHz spectral mask.
Figure 8: Indoor Spectral Compliance: BC-TAS Transmit Spectrum under IEEE 802.11n Mask.
The normalized Power Spectral Density (PSD) in Fig. 8 confirms that BC-TAS is fully compliant with the 802.11n mask even at low IBO levels. Sideband emissions are suppressed below 45-45 dB, allowing the transmitter to operate closer to the PA’s saturation region to maximize harvestable power for backscatter tags.
V-J Joint Optimization and Fundamental Trade-off
The flexibility of the BC-TAS framework is showcased through the fundamental trade-off between communication reliability (Primary Link BER) and the interference suppression benefit ($\Delta I$) afforded to the victim receiver.
Figure 9: Fundamental Pareto trade-off between primary link reliability and victim interference suppression in TGn Model F.
As illustrated in Fig. 9, the system exhibits a clear operating boundary with a distinct ”knee” that defines the limits of selection diversity. The smoothed boundary (Savitzky-Golay filtered) reveals the optimal operating ”knee” where significant suppression is achieved with minimal BER penalty. By increasing the victim weight $\lambda_{V}$, the system achieves a maximum interference suppression of approximately $6.5$ dB. The ”Optimal Operating Point” in the figure is identified at $\Delta I\approx 4.5$ dB, where the system achieves a significant $2.8\times$ reduction in linear interference power while maintaining a primary link BER on the order of $10^{-3}$.
This point represents the threshold of maximum marginal utility. In the region prior to this point ($0\leq\Delta I\leq 4.5$ dB), the BC-TAS algorithm successfully identifies antenna-subcarrier mappings that are naturally favorable to the victim with negligible impact on the primary link. However, beyond this $4.5$ dB threshold, the system enters a regime of diminishing returns; further gains in interference mitigation force the selection of antennas with significantly lower channel gains for the primary receiver, resulting in an exponential increase in BER as the diversity gain is exhausted to satisfy the strict coexistence constraint.
Figure 10: Pareto front scaling with antenna count $N_{tx}\in\{2,4,8\}$.
The impact of hardware complexity on this trade-off is further explored in Fig. 10. As $N_{tx}$ increases, the Pareto front shifts the entire operating boundary toward the high-suppression, low-BER region.. This shift indicates that higher-order selection diversity provides a more favorable trade-off surface, allowing for simultaneous improvements in both reliability and coexistence. For $N_{tx}=8$, the system can achieve suppression levels exceeding $7$ dB while remaining well below the BER levels of a $N_{tx}=2$ system with zero suppression.
V-K Spatial Correlation, PAPR, and Backscatter Efficiency
Conventional multi-user MIMO (MU-MIMO) systems and Space-Time Block Coding (STBC) schemes face significant challenges in WSN deployments, particularly regarding PAPR and backscatter energy harvesting efficiency. While linear receivers like MMSE detectors experience severe noise enhancement under high transmit-side spatial correlation ($\mathbf{R}_{\mathrm{tx}}$), STBC requires multiple active RF chains, which increases hardware cost and limits the power available for backscatter modulation.
In contrast, the proposed BC-TAS scheme addresses these bottlenecks by exploiting selection diversity across spatial and frequency domains. By activating only a single transmit antenna per subcarrier, BC-TAS inherently solves the high PAPR issue typically found in multi-stream MIMO, as the localized power concentration reduces the dynamic range of the transmitted signal. Furthermore, by identifying the most favorable propagation paths that maximize backscatter reflection power while minimizing interference, the scheme ensures high energy efficiency for passive sensor nodes.
Figure 11: Reliability robustness vs. transmit antenna correlation $\mathbf{R}_{\mathrm{tx}}$ at $\mathrm{SNR}=16$ dB.
As illustrated in Fig. 11, at a fixed SNR of $16$ dB, the MMSE baseline exhibits an exponential rise in BER, reaching approximately $2.8\times 10^{-2}$ at $\mathbf{R}_{\mathrm{tx}}=0.9$. While STBC provides better stability through orthogonal diversity, it maintains a performance gap significantly wider than the proposed scheme.
The proposed BC-TAS demonstrates superior robustness, achieving a BER of approximately $4.5\times 10^{-4}$ at extreme correlation levels ($0.9$). This result confirms that BC-TAS effectively circumvents the ill-conditioned channel matrices that plague MMSE and the hardware overhead of STBC. By focusing transmit energy into a single optimal antenna-subcarrier mapping, BC-TAS maximizes the signal-to-noise ratio for both the primary receiver and the backscatter tag, providing a low-complexity solution for high-density, correlated WSN environments.
V-L Physical Layer Insights: Frequency-Selective Notching
To provide physical intuition into the BC-TAS mechanism, we analyze the Spectral Power Density (SPD) of the transmitted signal as observed at the victim receiver in a frequency-selective TGn Model F environment.
Figure 12: SPD Analysis: Frequency-selective notching at the victim receiver.
As illustrated in Fig. 12, when the system operates in ”Protective Mode” ($\lambda_{V}=10$), the algorithm selectively maps subcarriers to antennas that exhibit deep fading at the victim’s spatial location. This results in a ”notching” effect where the interference power is surgically removed from specific subcarriers. For $\lambda_{V}=10$, the BC-TAS algorithm identifies and exploits natural channel nulls across the subcarrier index to protect the victim.
Crucially, this observed ”notching” validates the spatial-frequency decoupling inherent in our algorithm. It proves that the interference nulling is achieved through channel-aware selection diversity rather than expensive digital filtering or active cancellation hardware—computational overheads that a passive backscatter tag cannot support. By exploiting the natural multipath environment to ”hide” the signal from the victim, the BC-TAS framework justifies its low hardware complexity while maintaining high primary link power on subcarriers where the victim is naturally shielded.
VI Discussion and Theoretical Validation
In this section, we synthesize our numerical findings with the analytical framework derived in Section III and Appendix A. The alignment between the observed gains and the order-statistics model provides a robust validation of the BC-TAS framework.
VI-A Linking Field Flattening to Lemma 1
The results in Fig. 4 provide empirical confirmation of the ”Field Flattening” effect described in Lemma 1. The $1.1$ dB BCF advantage over OES at $N_{t}=32$ demonstrates that selection diversity can be dual-purposed for both link reliability and peak power suppression.
The use of 1,000 Monte Carlo trials confirms that this is not a transient benefit but a persistent statistical property of the BC-TAS heuristic. By smoothing the measurement noise through the Kalman update (Section III-C), the system effectively ”regularizes” the cost function, ensuring that the selected antenna-subcarrier mapping represents a stable propagation path rather than a noisy outlier. This stability is critical for low-complexity backscatter tags that lack the processing power to track high-variance reflected waveforms.
VI-B Linking Interference Suppression to SDR Expansion
The observed $2.4$ dB PAPR reduction and the corresponding interference suppression ”knee” in Fig. 9 directly validate the result in Lemma 1. As established analytically, the selection of the minimum interference path among $N_{t}=4$ antennas yields an expected power reduction scaling by $1/N_{t}$.
In the logarithmic domain, this theoretical shift is approximately $10\log_{10}(4)\approx 6.02$ dB. Our empirical maximum suppression of $6.5$ dB (see Fig. 9) aligns closely with this bound. The expansion of the SDR is therefore a direct consequence of this ”PDF squeezing” effect, where the background interference floor $\mathbb{E}[P_{B}|\lambda_{T}]$ is lowered, allowing the SAW sensor to operate in a significantly cleaner spectral environment.
VI-C Validation of the three-order-of-magnitude Outage Reduction
The massive reliability improvement reported in Table III is mathematically consistent with the diversity order $L=N_{t}$. Theoretically, the outage probability for selection diversity follows $P_{out}\approx(P_{out,SISO})^{N_{t}}$.
Given our SISO baseline outage of $0.105$ (at $16$ dB), the theoretical prediction for $N_{t}=4$ is $(0.105)^{4}\approx 1.2\times 10^{-4}$. Our simulated BC-TAS outage probability of $1.75\times 10^{-4}$ tracks this theoretical floor with high precision. The slight delta is attributed to the ”Selection Penalty” $\Delta$, where the system occasionally sacrifices the absolute maximum SNR to satisfy the weight constraints $\lambda_{T}$ and $\lambda_{V}$, as derived in Appendix A.C.
VI-D The PAPR-Linearity-Efficiency Loop
The $2.4$ dB gain in the PAPR CCDF (Fig. 6) serves as the hardware-level justification for the EVM and Energy Efficiency gains. By selecting antenna mappings that truncate the high-power tail of the OFDM distribution, BC-TAS forces the signal into the linear region of the Rapp PA model.
This enables the transmitter to reduce the IBO by approximately $3$ dB while maintaining 16-QAM compliance (see Fig. 5). This ”liberated” power is redirected toward tag illumination, increasing the harvesting efficiency $\eta_{H}$. Consequently, the system closes a performance loop: selection diversity reduces PAPR, which improves PA efficiency, which ultimately expands the power budget for the passive backscatter link.
VII Conclusion
This paper presented the BC-TAS MD-OFDM framework, a hardware-efficient approach to integrating passive SAW backscatter sensors into multi-antenna wireless networks. By shifting the selection logic from a purely SNR-centric approach to a multi-objective sensing-aware heuristic, we successfully addressed the critical bottleneck of primary-secondary link coexistence. The proposed BC-TAS achieves a $604.7\times$ reduction in link outage probability compared to SISO baselines at 16 dB SNR, a result mathematically consistent with the diversity order $L=N_{t}$ derived in our order-statistics model. Furthermore, the system maintains a BER of approximately $3.2\times 10^{-4}$ at 10 dB SNR, significantly outperforming industry-standard norm-based selection methods while remaining robust to severe CSI imperfections of up to $20\%$.
Beyond link reliability, the framework demonstrates significant hardware advantages. By activating only a single RF chain per subcarrier, BC-TAS achieves a $2.4$ dB reduction in the PAPR tail compared to conventional MU-MIMO baselines. This statistical advantage allows the transmitter to operate with approximately $3$ dB less input back-off while remaining compliant with IEEE 802.11n spectral masks, leading to a $25\times$ improvement in system energy efficiency. This hardware linearity is fundamentally linked to our coexistence optimization, where we identified a Pareto “knee” at $4.5$ dB of interference suppression. In this regime, frequency-selective notching expands the SDR by surgically lowering the background interference floor without compromising primary link integrity.
Ultimately, the BC-TAS framework provides a scalable, low-complexity solution for next-generation backscatter-assisted WSNs. By exploiting the natural multipath environment to shape the spatial-frequency interference profile, the system ensures high-fidelity communication and robust sensing coexistence in dispersive indoor environments. Future work will investigate the integration of this selection logic into massive MIMO arrays and its impact on multi-tag collision resolution.
Appendix A Derivation of Selection Penalty and Interference Profile
To derive the selection penalty $\Delta$ and the interference suppression gain, we analyze the order statistics of $N_{t}$ independent antenna channels. Let $G_{j}=\|\mathbf{h}_{j,k}^{(L)}\|^{2}$, $P_{T,j}=\|\mathbf{h}_{j,k}^{(T)}\|^{2}$, and $P_{V,j}=\|\mathbf{h}_{j,k}^{(V)}\|^{2}$ represent the channel power gains for the $j$-th antenna, which are independent and identically distributed (i.i.d.) exponential random variables with means $\sigma_{L}^{2}$, $\sigma_{T}^{2}$, and $\sigma_{V}^{2}$, respectively.
A-A Standard TAS Baseline
In a conventional Max-SNR TAS system, the selected gain $G_{\max}=\max\{G_{1},\dots,G_{N_{t}}\}$ follows the harmonic expectation:
$\mathbb{E}[G_{\max}]=\sigma_{L}^{2}\sum_{n=1}^{N_{t}}\frac{1}{n}\approx\sigma_{L}^{2}\ln(N_{t})+\gamma_{e}$ (13)
where $\gamma_{e}$ is the Euler-Mascheroni constant. This serves as the theoretical upper bound for primary link performance.
A-B Interference Suppression Scaling
The BC-TAS selection rule minimizes $C_{j}=\frac{1}{G_{j}}+\lambda_{T}P_{T,j}+\lambda_{V}P_{V,j}$. In the interference-limited regime ($\lambda_{T},\lambda_{V}\to\infty$), the selection logic effectively identifies the minimum of $N_{t}$ realizations. Let $Y=\min\{P_{T,1},\dots,P_{T,N_{t}}\}$. The CDF of the suppressed backscatter power $Y$ is:
$F_{Y}(y)=1-\left[1-F_{P_{T}}(y)\right]^{N_{t}}=1-e^{-\frac{N_{t}y}{\sigma_{T}^{2}}}$ (14)
Consequently, the probability of an outage at the tag, defined as $\mathcal{P}(P_{T}>\gamma)$, is given by $e^{-\frac{N_{t}\gamma}{\sigma_{T}^{2}}}$, which decreases exponentially as $N_{t}$ increases. This provides the mathematical justification for the massive gain reduction in the interference floor observed in the numerical analysis.
A-C The Selection Penalty $\Delta$
The selection penalty $\Delta$ quantifies the trade-off between primary link SNR and sensing stability. Since the BC-TAS index $j_{k}^{\star}$ is determined by the interference terms as $\lambda_{T},\lambda_{V}\to\infty$, the selected gain $G_{j_{k}^{\star}}$ becomes a random sample from the exponential distribution, independent of the maximum. Thus, the effective gain is simply $\mathbb{E}[G_{j_{k}^{\star}}]=\sigma_{L}^{2}$. We define the penalty $\Delta$ as:
$\Delta=\frac{\mathbb{E}[G_{j_{k}^{*}}]}{\mathbb{E}[G_{\max}]}=\left(\sum_{n=1}^{N_{t}}\frac{1}{n}\right)^{-1}=\frac{1}{H_{N_{t}}}$ (15)
where $H_{N_{t}}=\sum_{n=1}^{N_{t}}1/n$ is the $N_{t}$-th harmonic number. For $N_{t}=8$, $\Delta\approx 0.368$, implying a $4.34$ dB loss in primary SNR.
Remark: In practical regimes where $\lambda_{T}$ and $\lambda_{V}$ are finite, the achieved gain lies in the interval $\sigma_{L}^{2},\sigma_{L}^{2}H_{N_{t}}]$, allowing for a tunable trade-off via the MOFS cost function. This explains why simulation results under optimized weights often exhibit lower penalties than the ”worst-case” asymptotic limit derived here.

### v2

@@
 To ensure robustness in low-power environments, we integrate a Kalman-based channel smoothing layer that maintains selection stability even under severe Channel State Information (CSI) estimation errors of up to 20%.
-Numerical evaluations using the non-linear Rapp power amplifier model and IEEE 802.11be dispersive channel models demonstrate that BC-TAS achieves a nearly three-order-of-magnitude reduction in outage probability and a $25\times$ improvement in network energy efficiency compared to standard MU-MIMO baselines.
+Numerical evaluations using the non-linear Rapp power amplifier model and IEEE 802.11be dispersive channel models demonstrate that BC-TAS achieves a nearly three-order-of-magnitude reduction in outage probability and a $25\times$ improvement in network energy efficiency compared to standard Multi-User Multiple Input, Multiple Output (MU-MIMO) baselines.
 Furthermore, the proposed dual-envelope control mechanism achieves a $2.4$ dB reduction in the transmitter’s Peak-to-Average Power Ratio (PAPR), allowing for a $3$ dB reduction in required input back-off (IBO) while ensuring spectral mask compliance.
@@
 I Introduction
-The rapid proliferation of Internet of Things (IoT) devices and their integration into large-scale Wireless Sensor Networks (WSNs) has intensified the demand for hybrid communication architectures.
-Modern industrial and medical environments now require a single active hub—the illuminator—to support high-throughput data transmission to legitimate users while simultaneously energizing and interrogating passive, backscatter-based sensors [Zhang2025_SR, Wu2024_WPCN].
-However, this integration creates a fundamental physical-layer conflict: the wideband, high-power waveforms required for reliable primary communication often saturate the sensitive dynamic range of passive tags or cause prohibitive interference to coexisting victim receivers in the same spectral vicinity [Tan2025_Coexist].
-Conventional Multiple-Input Multiple-Output (MIMO) techniques, such as Minimum Mean Square Error (MMSE) precoding, are typically optimized for spectral efficiency.
-However, they are increasingly unsuitable for the next generation of energy-constrained Passive WSNs (PWSNs) due to three critical limitations.
-First, the requirement for multiple active RF chains imposes a hardware complexity and power consumption floor that is incompatible with green-WSN goals [Zhao2024_Green].
-Second, the high Peak-to-Average Power Ratio (PAPR) inherent in multi-stream Orthogonal Frequency Division Multiplexing (OFDM) necessitates significant Power Amplifier (PA) back-off, leading to spectral regrowth and accelerated battery depletion [Park2024_PAPR].
-Most importantly, current precoding strategies lack a mechanism to regulate the incident RF envelope perceived by remote passive devices [Wang2024_Sense].
-For time-modulated sensors, such as Surface Acoustic Wave (SAW) tags, high-PAPR fluctuations act as a source of “envelope noise,” masking low-power acoustic reflections and degrading sensing fidelity [Liu2025_SAW].
-While diversity-oriented selection methods, such as Transmit Antenna Selection (TAS) and Multi-Dimensional OFDM (MD-OFDM), provide a low-complexity alternative with inherent PAPR benefits, they remain largely “sensing-blind” [Garcia2024_IM].
-Existing selection literature focuses exclusively on maximizing desired-link Signal-to-Noise Ratio (SNR).
-Consequently, there is an urgent need for a framework that treats the spatial-frequency resource blocks of an OFDM symbol as a tool for multi-objective field shaping—simultaneously optimizing primary link reliability, stabilizing backscatter energy incidence, and surgically nulling interference toward victim nodes [Ibrahim2025_MultiObj].
+The proliferation of Internet of Things (IoT) devices in large-scale Wireless Sensor Networks (WSNs) has intensified the demand for hybrid communication architectures.
+Modern industrial and medical environments require a single active hub—the illuminator—to support high-throughput data transmission while simultaneously energizing and interrogating passive, backscatter-based sensors [Zhang2025_SR, Wu2024_WPCN].
+However, this integration creates a fundamental physical-layer conflict: the wideband, high-power waveforms required for primary communication often saturate the sensitive dynamic range of passive tags or cause prohibitive interference to coexisting victim receivers [Tan2025_Coexist].
+Conventional Multiple-Input Multiple-Output (MIMO) techniques, such as Minimum Mean Square Error (MMSE) precoding, are increasingly unsuitable for energy-constrained Passive WSNs (PWSNs) due to three critical limitations.
+First, multiple active RF chains impose a hardware complexity and power floor incompatible with green-WSN goals [Zhao2024_Green].
+Second, the high Peak-to-Average Power Ratio (PAPR) of multi-stream OFDM necessitates significant Power Amplifier (PA) back-off, leading to spectral regrowth and accelerated battery depletion [Park2024_PAPR].
+Most importantly, current strategies lack a mechanism to regulate the incident RF envelope perceived by remote passive devices [Wang2024_Sense].
+For time-modulated sensors, such as Surface Acoustic Wave (SAW) tags, high-PAPR fluctuations act as “envelope noise,” masking low-power acoustic reflections and degrading sensing fidelity [Liu2025_SAW].
+While selection methods like Transmit Antenna Selection (TAS) and Multi-Dimensional OFDM (MD-OFDM) provide low-complexity alternatives with inherent PAPR benefits, they remain largely “sensing-blind” [Garcia2024_IM].
+Existing literature focuses almost exclusively on maximizing desired-link Signal-to-Noise Ratio (SNR).
+Consequently, there is an urgent need for a framework that treats spatial-frequency resource blocks as tools for multi-objective field shaping—simultaneously optimizing primary link reliability, stabilizing backscatter energy incidence, and surgically nulling interference toward victim nodes [Ibrahim2025_MultiObj].
 To address these gaps, we propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework.
-This is the first per-subcarrier selection scheme designed to bridge the “Illuminator’s Gap” by actively managing the RF conditions perceived by passive devices.
-By shifting the selection logic from a purely SNR-centric approach to a multi-objective sensing-aware heuristic, BC-TAS transforms the illuminator to act as a network-wide controller of both signal quality and incident energy stability.
-The central innovation of this work is the development of a unified cost function that embeds three interacting channel components—the legitimate link, tag-side incidence, and victim-link interference—into a per-subcarrier decision rule.
-Furthermore, we leverage the sparsity of MD-OFDM to achieve a dual-regulation effect: minimizing transmitter-side PAPR to ensure PA linearity while simultaneously minimizing the Backscatter Crest Factor (BCF) to ensure tag-side envelope stability.
-This dual-regulation ensures that the system remains compliant with strict spectral masks, such as those defined in the maturing IEEE 802.11 Ambient Power Communication (AMP) frameworks, even when operating in highly dispersive environments [IEEE80211AMP_Update, IEEE80211bk_2025].
+To the best of our knowledge, this is the first per-subcarrier selection scheme designed to bridge the “Illuminator’s Gap” by actively managing the RF conditions perceived by passive devices.
+By shifting selection logic from an SNR-centric approach to a multi-objective sensing-aware heuristic, BC-TAS transforms the illuminator into a network-wide controller of both signal quality and incident energy stability.
+The central innovation is a unified cost function embedding three interacting channel components—the legitimate link, tag-side incidence, and victim interference—into a per-subcarrier decision rule.
+We leverage the sparsity of MD-OFDM to achieve a dual-regulation effect: minimizing transmitter-side PAPR for PA linearity while simultaneously minimizing the Backscatter Crest Factor (BCF) for tag-side envelope stability.
+This ensures compliance with strict spectral masks, such as the IEEE 802.11bp Ambient Power Communication (AMP) frameworks, even in highly dispersive environments [IEEE80211AMP_Update].
 The key contributions of this work are summarized as follows:
-  • BC-TAS Framework Design: We introduce a novel multi-objective per-subcarrier selection metric that explicitly balances primary link reliability against the stability of backscatter incidence and the suppression of victim-link interference.
-  • Analytical Metric for Sensing Fidelity (BCF): We define and derive the BCF as a novel performance metric to quantify the temporal stability of the RF envelope at the tag.
-We provide a theoretical bound for BCF reduction under $N_{t}$-order selection diversity, establishing the first formal link between spatial resource allocation and sensing ”envelope noise.
-  • Dual-Envelope Regulation: We establish a joint control mechanism that exploits sparse subcarrier activation to simultaneously reduce the transmitter’s PAPR and the tag’s incident BCF, providing the first quantitative link between illuminator antenna selection and backscatter sensing fidelity.
-  • Low-Complexity Heuristic: We formulate the BC-TAS decision as a constrained optimization problem and develop a low-complexity greedy solution incorporating recursive Kalman-based channel smoothing to ensure selection stability under imperfect CSI.
-The algorithm scales linearly with the number of subcarriers, providing a 2.3× execution speedup over traditional MMSE detectors while maintaining robustness to rapid channel fluctuations.
-  • Non-linear Hardware Validation: We perform a comprehensive evaluation using the Rapp PA model and IEEE 802.11n/be channel models, demonstrating over two orders of reduction in outage probability and a $25\times$ improvement in system energy efficiency compared to standard SISO and MIMO baselines.
+  • BC-TAS Framework Design: A novel multi-objective per-subcarrier selection metric that balances primary link reliability against backscatter incidence stability and victim interference suppression.
+  • Analytical Metric for Sensing Fidelity BCF: We define the BCF to quantify the temporal stability of the RF envelope at the tag and provide a theoretical bound for BCF reduction under $N_{t}$-order selection diversity.
+  • Dual-Envelope Regulation: A joint control mechanism exploiting sparse subcarrier activation to simultaneously reduce transmitter PAPR and tag incident BCF, linking antenna selection directly to sensing fidelity.
+[…]
+  • Hardware Aware Simulation: Comprehensive evaluation using Rapp PA and IEEE 802.11 TGn models, demonstrating two orders of magnitude reduction in outage probability and a $25\times$ improvement in energy efficiency.
 The remainder of this paper is organized as follows.
-Section II reviews the state-of-the-art in TAS, backscatter sensing, and PAPR-aware transmission.
-Section III details the system model and the MOFS formulation, including the Kalman-based channel smoothing.
-Section IV presents the proposed greedy selection algorithm and derives its theoretical diversity and scaling properties.
-Section V provides extensive simulation results, including BER, BCF, and hardware efficiency metrics.
-Section VI synthesizes these findings with a theoretical validation of the observed gains and trade-offs.
-Finally, Section VII concludes the work, followed by the derivation of the selection penalty in Appendix A.
+Section II reviews the state-of-the-art.
+Section III details the system model and Multi-Objective Field Shaping (MOFS) formulation.
+Section IV presents the greedy selection algorithm and its theoretical properties.
+Section V provides extensive simulation results and synthesizes these with a theoretical validation.
+Section VI concludes the work.
 II Related Work
-II-A Transmit Antenna Selection (TAS)
-TAS remains a cornerstone for low-complexity diversity in modern wireless architectures.
-Recent research by Kumar et al.
-(2025) demonstrates that TAS provides significant diversity gains in active RIS-assisted downlink NOMA systems, where the diversity order scales jointly with BS antennas and RIS elements [Kumar2025].
-Similarly, Han et al.
-(2025) emphasize that TAS-based strategies, particularly under statistical Channel State Information (CSI), can effectively improve outage probability and sum throughput without the prohibitive RF-chain costs of full MIMO [Han2025].
-However, these classical and 6G-era TAS frameworks primarily optimize internal communication metrics (SNR, Capacity, BER).
-There remains a critical research gap in applying TAS to manage external physical-layer constraints, such as the RF envelope of a backscatter tag or the interference profile of a sensitive victim node.
-II-B OFDM PAPR Reduction and Hardware Linearity
+II-A Transmit Antenna Selection (TAS) and Index Modulation
+TAS remains a cornerstone for low-complexity diversity.
+Recent research demonstrates significant gains in RIS-assisted NOMA systems and 6G-era architectures where diversity scales with antenna count without prohibitive RF-chain costs [Kumar2025, Han2025].
+Parallelly, Index Modulation (IM) and MD-OFDM have emerged as massive connectivity enablers by exploiting subcarrier indices to embed information [Li2023].
+While IM provides sparse activation and reduced interference [Mgobhozi2023, VilaInsa2025], existing TAS and IM literature primarily optimizes internal metrics like SNR and Bit Error Rate (BER).
+A critical gap remains in utilizing these selection variables to manage external constraints, such as the RF envelope of a backscatter tag.
+II-B PAPR, Hardware Linearity, and Energy Efficiency
 PAPR is the primary driver of PA non-linearity, leading to spectral regrowth and constellation distortion.
-Contemporary work in Integrated Sensing and Communication (ISAC) has highlighted that PAPR reduction is essential for improving transmitter efficiency and range under non-linear High Power Amplifier (HPA) constraints [Huang2022].
-While probabilistic methods like Partial Transmit Sequences (PTS) have been refined to embed side information implicitly to reduce overhead [Goel2022], they still introduce signal processing latency.
-From a system-level perspective, Komala (2024) demonstrates that inefficient transmission directly translates to reduced network longevity in WSNs [Komala2024].
-By mapping these findings to the Rapp model, it becomes evident that an inherently low-PAPR waveform, rather than an algorithmic mitigation layer, is vital for long-lived, energy-efficient Wireless Power and Data Transfer (WPDT).
-II-C Index Modulation and MD-OFDM
-Index Modulation (IM) and its variants, such as MD-OFDM, have emerged as core 6G enablers for massive connectivity [Li2023].
-Unlike classical OFDM, IM exploits the indices of subcarriers or antennas to embed information, providing a sparse activation pattern that can lead to lower PAPR and reduced inter-channel interference [Mgobhozi2023].
-Recent advances in permutational IM have even enabled low-complexity, non-coherent detection that is robust under antenna correlation [VilaInsa2025].
-However, most IM literature focuses on maximizing spectral efficiency (SE) or reliability (BER).
-Even when PAPR is addressed, as in multilevel dither signal designs for OFDM-IM [Kim2018], the focus is purely on internal signal integrity.
-Our work bridges this gap by treating the MD-OFDM index selection as a first-class design variable for shaping the incident energy at a backscatter sensor.
-II-D Backscatter Communications and Symbiotic Radio
-The evolution of backscatter has moved toward multi-antenna and broadband configurations to improve link reliability in dispersive channels [Zhang2017, Chen2024].
-Optimization efforts have traditionally been split: either optimizing tag reflection coefficients for energy-rate trade-offs [Goay2024] or utilizing deep reinforcement learning for beamforming in bistatic setups [Zargari2024].
-Recent symbiotic radio research has proposed strategic subcarrier nulling to eliminate direct-link interference in OFDM backscatter [Janjua2025].
-Despite these advances, the ”Illuminator’s Gap” persists; few studies investigate using the transmitter’s resource allocation (like BC-TAS) to actively shape the temporal stability of the RF envelope, which is critical for time-modulated sensors like SAW tags.
-II-E Coexistence in Passive SAW-based Sensor Networks
-As standardization bodies like IEEE 802.11 extend positioning to 320 MHz channels [IEEE80211bk], the need for interference-resilient signaling becomes paramount.
-Passive SAW sensors represent a unique challenge in this landscape, as they operate without batteries by converting incident RF energy into time-delayed acoustic reflections [Chen2024, Zhang2017].
-Unlike active IoT devices, PWSNs are highly susceptible to:
-  • RF Envelope Instability: High PAPR signals create power fluctuations that mask the low-power acoustic reflections [Huang2022, Kim2018].
-  • Spectral Competition: While frequency-domain nulling has been proposed [Liu2025, Janjua2025], these methods often ignore the hardware non-linearities of the illuminator.
-Industry consensus in the IEEE 802.11 AMP group highlights a critical gap: the absence of selection-based mechanisms that utilize the transmitter’s spatial degrees of freedom to optimize the energy envelope at the tag while protecting primary link reliability.
-The performance of selection-based architectures in industrial settings is inextricably linked to the underlying radio propagation environment.
-Prior work has introduced Variational Autoencoder (VAE) based models to predict signal-to-interference-plus-noise ratio (SINR) heatmaps within automated Industry 4.0 warehouses, demonstrating how interacting objects within smart infrastructure significantly impact 5G propagation [GuliaVAE2025].
-Furthermore, detailed evaluations of 60 GHz millimeter-wave connectivity in these environments reveal that non-line-of-sight (nLOS) paths created by metallic shelving and autonomous handling agents introduce complex multipath fading effects [Gulia60GHz2023].
-To manage these dynamics without high computational costs, stabilized symbolic regression has been utilized to derive compact, interpretable expressions for block error rate (BLER) prediction, offering neural-network-level accuracy suitable for real-time embedded deployment [GuliaSymbolic2024].
-These studies underscore the necessity for low-complexity, environment-aware selection mechanisms like BC-TAS that can adapt to the complex spatial profiles of automated warehouses.
-Recognizing these coexistence challenges, the industrial shift toward sustainable IoT is evidenced by the formation of the IEEE 802.11 AMP Task Group (TGbp).
-According to the project authorization scope, the group aims to enable battery-free station (STA) operation through energy harvesting and ultra-low-complexity signaling [IEEE80211AMP_PAR].
-A primary challenge identified in the group’s functional requirements is the coexistence of these passive-like devices with legacy high-throughput STAs [IEEE80211AMP_Update].
-While the task group explores physical-layer modifications for 2.4 GHz and sub-1 GHz bands, current proposals primarily focus on frequency-domain resource allocation.
-Our work, BC-TAS, complements these efforts by introducing a spatial-domain selection mechanism that addresses the hardware non-linearities and peak-envelope requirements that are not yet fully specified in the maturing AMP framework.
+In Integrated Sensing and Communication (ISAC), PAPR reduction is vital for efficiency under non-linear High Power Amplifier (HPA) constraints [Huang2022].
+While probabilistic methods like Partial Transmit Sequences (PTS) reduce PAPR, they introduce processing latency that impacts WSN longevity [Goel2022, Komala2024].
+Mapping these constraints to the Rapp model suggests that inherently low-PAPR waveforms are superior to algorithmic mitigation for energy-efficient Wireless Power and Data Transfer (WPDT).
+II-C Backscatter Communications and Symbiotic Radio
+Backscatter evolution has moved toward multi-antenna and broadband configurations to enhance link reliability [Chen2024].
+Optimization efforts often split between tag-side reflection coefficients and bistatic beamforming [Goay2024, Zargari2024].
+While symbiotic radio research proposes subcarrier nulling to eliminate direct-link interference [Janjua2025], few studies investigate using transmitter resource allocation to shape the temporal stability of the RF envelope—a factor critical for time-modulated SAW tags.
+II-D Coexistence in Passive SAW-based Sensor Networks
+As IEEE 802.11 positioning extends to 320 MHz channels [IEEE80211bk], interference-resilient signaling is paramount.
+Passive SAW sensors are particularly susceptible to RF envelope instability, where high PAPR masks low-power acoustic reflections [Huang2022, Kim2018].
+While frequency-domain nulling has been proposed [Liu2025, Janjua2025], these methods often ignore hardware non-linearities.
+Industry consensus within the IEEE 802.11 AMP group (TGbp) highlights the lack of selection-based mechanisms to optimize tag energy envelopes while protecting primary link reliability [IEEE80211AMP_Update].
+Selection performance in industrial settings is inextricably linked to the propagation environment.
+Prior work utilizing VAE-based SINR heatmaps and 60 GHz connectivity analysis in automated warehouses reveals complex multipath fading caused by metallic infrastructure [GuliaVAE2025, Gulia60GHz2023].
+To manage these dynamics, interpretable symbolic regression has been used for real-time BLER prediction [GuliaSymbolic2024].
+These studies underscore the necessity for environment-aware mechanisms like BC-TAS.
+Our framework complements current TGbp efforts by introducing a spatial-domain selection mechanism that addresses hardware non-linearities and peak-envelope requirements not yet fully specified in the maturing AMP framework [IEEE80211AMP_PAR].
 III System Model and BC-TAS Formulation
@@
 As shown in Section V, the performance gap between the greedy BC-TAS and the exhaustive search remains below $0.5$ dB in the high-SNR regime, justifying the use of the low-complexity heuristic for practical deployments.
-V Simulation and Results
-This section details the simulation environment, defines the key parameters used, and presents the performance metrics of the proposed Backscattered Constrained Transmit Antenna Selection Multi-Dimensional Orthogonal Frequency Division Multiplexing (BC-TAS MD-OFDM) model compared to baseline schemes.
-V-A Simulation Setup
-The system performance was evaluated using Monte Carlo simulations.
-The core parameters are summarized in Table I.
-To ensure relevance to emerging standards, the channel environment is modeled based on the IEEE 802.11be (Wi-Fi 7) TGbe indoor specifications, utilizing Multipath Fading Models (TGn A–F) to represent varying degrees of frequency selectivity in warehouse environments.
-To characterize hardware-induced distortions, we incorporate a non-linear Rapp PA model at the transmitter.
-The AM/AM conversion follows the smoothness factor $p=2$, which represents a typical Solid State Power Amplifier (SSPA).
-This model is critical for evaluating the impact of high-PAPR and high-BCF waveforms on spectral regrowth and Legitimate-link Bit Error Rate (BER).
-Simulations were performed on a high-performance computing node equipped with an NVIDIA Tesla V100-PCIE-16GB GPU and Intel Xeon processors, utilizing the NVIDIA CUDA 12.3 compiler driver.
-Execution times reported in Table II represent the total duration required to process $10^{7}$ bits across an SNR range of 0 to 28 dB.
+V Computational Complexity Analysis
+The feasibility of BC-TAS for real-time industrial deployment is rooted in its linear scaling properties.
+Unlike conventional MMSE precoding, which requires matrix inversions of the form $\mathcal{O}(K^{3})$ (where $K$ is the number of users/streams), the proposed greedy selection algorithm operates on a per-subcarrier basis with a complexity of $\mathcal{O}(N_{SC}\cdot N_{t})$.
+Specifically, for each subcarrier $n\in\{1,\dots,N_{SC}\}$, the metric $M(n,i)$ is evaluated $N_{t}$ times.
+The inclusion of the scalar Kalman Filter adds a constant overhead of $\mathcal{O}(1)$ per subcarrier, as it avoids high-dimensional state-space computations.
+As a result, for a standard 1024-subcarrier system with 4 transmit antennas, BC-TAS reduces the total floating-point operations (FLOPs) by approximately $64\%$ compared to full-complexity MIMO baselines.
+This reduction directly translates to the $2.3\times$ execution speedup reported in our results, enabling the illuminator to adapt to the fast-fading environments of automated warehouses.
+VI Simulation and Results
+This section evaluates the proposed BC-TAS MD-OFDM framework against state-of-the-art (SOTA) baselines across indoor dispersive environments.
+VI-A Simulation Setup
+System performance is evaluated via Monte Carlo simulations using the parameters in Table I.
+To ensure standard relevance, the channel is modeled on IEEE 802.11 TGn specifications, utilizing Multipath Fading Models (TGn A–F) to represent varying frequency selectivity in warehouse environments.
+Hardware distortions are characterized by a non-linear Rapp PA model ($p=2$) at the transmitter to evaluate the impact of high-PAPR/BCF waveforms on spectral regrowth and Legitimate-link Bit Error Rate (BER).
+Simulations were conducted on a high-performance node (NVIDIA Tesla V100 GPU, Intel Xeon processors) with the CUDA 12.3 driver.
+Execution times in Table II reflect the processing of $10^{7}$ bits across an SNR range of 0–28 dB.
 TABLE I: Key Simulation Parameters
-Parameter Category Parameter Value
-System and Channel Model
-System Scheme Proposed BC-TAS MD-OFDM
-Baseline Schemes Conventional TAS / MMSE $N_{T}=4,N_{R}=4$
-Standard Context IEEE 802.11be EHT Physical Layer
-Channel Model Indoor Multipath Fading TGn Models A–F
-Fading Distribution Rayleigh Independent Fading Paths
-CSI Assumption P-CSI / I-CSI Perfect / Estimation Error $\sigma_{e}^{2}$
-OFDM and PHY Layer
-Subcarrier Count ($N$) $N_{\mathrm{SC}}$ 256 / 1024
-Cyclic Prefix Length ($N_{\mathrm{CP}}$) $L_{\mathrm{CP}}$ 32 samples
-Modulation Scheme $M$-QAM 16-QAM / 64-QAM
-Total Simulated Bits $N_{\mathrm{bits}}$ $10^{7}$ bits
-Non-Linearity and Field Shaping
-PA Model Rapp Model $p=2$ (Smoothness factor)
-MOFS Weights $\lambda_{T},\lambda_{V}$ $[0.1,0.5,1.0]$
-Tag/Victim Distances $d_{L},d_{T},d_{V}$ 10 m, 2 m, 5 m
-Kalman Process Noise ($Q$) $Q_{KF}$ $1\times 10^{-4}$
-Kalman Meas. Noise ($R$) $R_{KF}$ $1\times 10^{-2}$
-Monte Carlo Trials $N_{trials}$ 1,000
-V-B Comparative Analysis with State-of-the-Art (SOTA) Baselines
-To validate the competitive standing of the proposed BC-TAS framework, we benchmarked its performance against four distinct SOTA categories: fixed-diversity baselines [SOTA_2_MU_STBC], low-complexity selection methods [SOTA_1_MMSE], theoretical upper bounds (OES) [SOTA_3_Exhaustive_Search], and backscatter-specific multi-antenna contexts [SOTA_4_Backscatter_Multi_Antenna].
+Parameter Value / Configuration
+System Scheme BC-TAS MD-OFDM ($N_{T},N_{R}=4$)
+Standard Context IEEE 802.11 TGn (EHT PHY)
+Channel Model TGn Models A–F (Rayleigh)
+CSI Assumption Perfect / $\sigma_{e}^{2}$ Error (5%, 20%)
+Subcarriers ($N_{\mathrm{SC}}$) 256 (80 MHz BW)
+Cyclic Prefix ($L_{\mathrm{CP}}$) 32 samples
+Modulation ($M$-QAM) 16-QAM / 64-QAM
+Simulated Bits ($N_{\mathrm{bits}}$) $10^{7}$ bits
+PA Model (Rapp) $p=2$, $V_{sat}=1$
+MOFS Weights $\lambda_{T},\lambda_{V}\in\{0.1,0.5,1.0\}$
+Distances ($d_{L},d_{T},d_{V}$) 10 m, 2 m, 5 m
+Kalman ($Q_{KF},R_{KF}$) $10^{-4},10^{-2}$
+Monte Carlo Trials 1,000
+VI-B Comparative Analysis with SOTA Baselines
+To validate the BC-TAS framework, we benchmark it against four SOTA categories: fixed-diversity baselines [SOTA_2_MU_STBC], low-complexity selection [SOTA_1_MMSE], theoretical upper bounds optimal exhaustive search (OES) [SOTA_3_Exhaustive_Search], and backscatter-specific multi-antenna configurations [SOTA_4_Backscatter_Multi_Antenna].
 Figure 1: BER comparison of the proposed BC-TAS framework against SOTA baselines in TGn Model F.
-As shown in the simulation results in Fig.
-1, the proposed BC-TAS significantly outperforms the industry-standard Norm-Based Antenna Selection (NBAS) [SOTA_5_NBAS].
-At an average SNR of 10 dB, BC-TAS achieves a BER of $\approx 3.2\times 10^{-4}$, whereas NBAS remains at $\approx 2.6\times 10^{-2}$.
-This order-of-magnitude improvement is attributed to our model’s ability to exploit frequency-selective diversity rather than relying on a static norm-based metric.
-The proposed scheme achieves a near-optimal slope while maintaining robustness to high CSI error ($20\%$).
-V-C Impact of CSI Imperfection and Robustness Analysis
-A critical requirement for practical backscatter deployments is resilience to CSI estimation errors, which are prevalent due to the low-power nature of the tag.
-We evaluate the robustness of the proposed BC-TAS framework under three regimes: Perfect CSI (P-CSI), 5% estimation error (csi_error_low=0.05csi\_error\_low=0.05), and a severe 20% error condition (csi_error_high=0.2csi\_error\_high=0.2).
+As shown in Fig.
+1, BC-TAS significantly outperforms the industry-standard Norm-Based Antenna Selection (NBAS) [SOTA_5_NBAS].
+At 10 dB SNR, BC-TAS achieves a BER of $\approx 3.2\times 10^{-4}$, while NBAS remains at $\approx 2.6\times 10^{-2}$.
+This order-of-magnitude improvement stems from the framework’s ability to exploit frequency-selective diversity across subcarriers rather than relying on a static, wideband norm-based metric.
+Furthermore, the proposed scheme achieves a near-optimal diversity slope and remains robust even under severe 20% CSI error conditions.
+VI-C Impact of CSI Imperfection and Robustness Analysis
+Practical backscatter deployments require resilience to CSI estimation errors, which are prevalent due to the low-power nature of the tag.
+We evaluate the BC-TAS framework under three regimes: Perfect CSI (P-CSI), 5% estimation error, and 20% error.
 As illustrated in Fig.
-1, the BC-TAS (Perfect) configuration defines the heuristic’s lower bound, tracking closely with the MD-OFDM (Max Gain) ceiling until approximately 10 dB SNR.
-In this regime, the system effectively exploits the full spatial-frequency diversity available in the TGn Model F environment.
-  • Moderate Imperfection (5% Error): Under a 5% CSI estimation error, the BC-TAS model exhibits a negligible SNR penalty of approximately 1.5 dB at a target BER of $10^{-3}$.
-Notably, the diversity slope remains steep, indicating that the multi-objective selection metric is still capable of identifying favorable subcarrier-antenna mappings despite minor phase and magnitude perturbations.
-  • Severe Imperfection (20% Error): At 20% error, the system encounters a performance floor at a BER of $\approx 1.8\times 10^{-2}$.
-While the high error rate significantly degrades the selection accuracy, the proposed BC-TAS (20% Error) remarkably maintains a lower BER than the Norm-Based TAS and MMSE 4x4 baselines in the low-to-mid SNR regions (0–8 dB).
-The distinct robustness of the proposed model over NBAS is attributed to the fundamental difference in their selection logic.
-While the performance of NBAS [SOTA_5_NBAS] collapses under severe 20% CSI error—often erroneously selecting a ”weak” antenna path due to noisy norm estimations—the multi-objective BC-TAS remains more resilient.
-By considering the average channel state across all 256 subcarriers through its heuristic metric, BC-TAS provides a ”smoothing” effect against noise.
-This prevents the selection of consistently poor paths, ensuring that the probability of selecting a deep-fade antenna across the entire bandwidth is significantly minimized, thereby maintaining a reliable communication link even under high CSI aging or estimation error.
-V-D Computational Efficiency vs. Performance Ceiling
-A critical finding of this study is the relationship between the proposed heuristic and the Optimal Exhaustive Search (OES) baseline.
-While OES [SOTA_3_Exhaustive_Search] defines the absolute performance ceiling—achieving zero BER at 8 dB—it requires a per-carrier exhaustive search across all $N_{sc}$ subcarriers.
-[…]
+1, the P-CSI configuration tracks the MD-OFDM ceiling until 10 dB SNR, effectively exploiting the full spatial-frequency diversity of the TGn Model F environment.
+Under moderate imperfection (5% error), BC-TAS exhibits a negligible 1.5 dB SNR penalty at a BER of $10^{-3}$ while maintaining a steep diversity slope.
+Under severe imperfection (20% error), the system encounters a performance floor at $\approx 1.8\times 10^{-2}$.
+Despite this, BC-TAS (20% Error) maintains a lower BER than both NBAS and MMSE $\times4$ baselines in the 0–8 dB SNR range.
+This robustness stems from the selection logic: while NBAS [SOTA_5_NBAS] collapses by erroneously selecting weak paths due to noisy norm estimations, the multi-objective BC-TAS provides a ”smoothing” effect.
+By considering the aggregate channel state across all 256 subcarriers, BC-TAS minimizes the probability of selecting deep-fade antennas, ensuring link reliability even under high CSI aging or estimation error.
+VI-D Computational Efficiency vs. Performance Ceiling
+While OES [SOTA_3_Exhaustive_Search] defines the absolute performance ceiling (zero BER at 8 dB), it requires a per-carrier exhaustive search across all $N_{sc}$ subcarriers.
+As shown in Table II, our BC-TAS heuristic demonstrates a $2.3\times$ speedup over conventional MMSE detectors, completing the simulation in 79.16 seconds compared to 181.69 seconds for MMSE $\times4$.
+In GPU-accelerated environments, BC-TAS achieves execution times comparable to OES while bridging the gap between theoretical optimality and practical real-time constraints.
+This latency reduction is vital for energy-constrained WSNs requiring rapid selection to maintain the sensing-communication link.
 TABLE II: Computational Complexity and Execution Time Benchmark
@@
 Proposed BC-TAS $\mathbf{\mathcal{O}(N_{T}N_{R})}$ Heuristic Multi-Objective 79.16
-V-E BER Performance and Environmental Robustness
-The primary performance metric is the BER versus the SNR. The diversity gain provided by the BC-TAS algorithm is demonstrated by the steeper slope of the BER curve compared to the MMSE baseline.
+VI-E BER Performance and Environmental Robustness
+The diversity gain of BC-TAS is evidenced by the significantly steeper BER slope compared to the MMSE baseline.
+To validate robustness, performance is analyzed across IEEE 802.11 TGn channel models A–F (Fig.
+2).
 Figure 2: BER performance of the proposed BC-TAS MD-OFDM across various IEEE 802.11TGn indoor environments (Models A–F).
-To validate the model’s robustness, performance is analyzed across the full suite of TGn channel models (A–F), as illustrated in Fig.
-2.
-We observe that while the system maintains high reliability in all scenarios, the environment significantly impacts the diversity order.
-Model A (Flat/Small Office) exhibits the best performance due to the lack of severe frequency selectivity.
-In contrast, as we move toward Model F (Large Open Space) with higher delay spreads, the increased frequency-selective fading introduces more frequent deep fades.
-However, the MD-OFDM selection diversity effectively mitigates these effects, achieving a $10^{-3}$ BER at approximately 12–14 dB SNR across all dispersive models.
-This demonstrates the framework’s ability to maintain a consistent sensing-communication link regardless of the physical indoor layout.
-V-F Outage Probability and Diversity Gain Analysis
-A critical bottleneck for WSNs is the prevalence of deep fading in indoor environments, which leads to link outages and energy-inefficient retransmissions.
-To evaluate the robustness of BC-TAS, we analyze the Cumulative Distribution Function (CDF) of the instantaneous SINR, $\gamma$, in a frequency-selective TGn Model D environment.
-The outage probability is defined as $P_{\mathrm{out}}=P(\gamma<\gamma_{\mathrm{th}})$, where the target threshold is set to $\gamma_{\mathrm{th}}=7$ dB.
-This threshold represents the minimum operating SINR required for reliable QPSK demodulation and successful packet decoding in standard low-power protocols.
-Figure 3: Outage Probability CDF in TGn Model D at an average $\mathrm{SNR}=16$ dB.
-The simulation results, illustrated in Fig.
-3 on a semi-logarithmic scale, reveal that the proposed BC-TAS scheme significantly reshapes the SINR distribution.
-While conventional random selection closely tracks the Theoretical Rayleigh (SISO) baseline—exhibiting an outage probability of approximately 0.10580.1058—the BC-TAS scheme suppresses the lower tail of the distribution through high-order selection diversity.
-The proposed BC-TAS scheme demonstrates a significantly steeper diversity slope compared to the Theoretical Rayleigh (SISO) and conventional random selection baselines.
-Specifically, at the $7$ dB threshold, BC-TAS achieves an outage probability of $0.000175$, representing a $604.77\times$ reduction in link outages.
+While Model A (Flat/Small Office) yields the best performance due to low frequency selectivity, the diversity order remains high even in Model F (Large Open Space).
+Despite the increased delay spreads and deep fades in larger environments, the MD-OFDM selection diversity achieves a $10^{-3}$ BER at approximately 12–14 dB SNR across all dispersive models.
+This confirms the framework’s capability to maintain consistent sensing-communication links regardless of the physical indoor layout.
+VI-F Outage Probability and Diversity Gain Analysis
+To evaluate link reliability in deep-fading indoor environments, we analyze the Cumulative Distribution Function (CDF) of the instantaneous SINR, $\gamma$, in a TGn Model F environment.
+The outage probability is defined as $P_{\mathrm{out}}=P(\gamma<\gamma_{\mathrm{th}})$, where $\gamma_{\mathrm{th}}=7$ dB represents the minimum threshold for reliable QPSK demodulation.
+Figure 3: Outage Probability CDF in TGn Model F at an average $\mathrm{SNR}=16$ dB.
+As illustrated in Fig.
+3, BC-TAS significantly reshapes the SINR distribution.
+While conventional random selection tracks the Theoretical Rayleigh (SISO) baseline ($P_{\mathrm{out}}\approx 0.1058$), BC-TAS suppresses the distribution’s lower tail.
+At the $7$ dB threshold, BC-TAS achieves an outage probability of $1.75\times 10^{-4}$, representing a $604.77\times$ reduction in link outages.
+As detailed in Table III, this improvement stems from the transition to a diversity order of $L=4$, where reliability gains grow exponentially with $N_{T}$.
+In Rayleigh fading, $P_{\mathrm{out}}$ scales as $(P_{\mathrm{out,SISO}})^{N_{T}}$; thus, our three-order-of-magnitude reduction aligns with theoretical expectations.
+By bypassing subcarriers in deep fades, BC-TAS ensures the signal remains above the decoding threshold $\approx 99.99\%$ of the time.
+Our simulated outage of $1.75\times 10^{-4}$ closely tracks the theoretical floor $(P_{\mathrm{out,SISO}})^{N_{t}}\approx 1.2\times 10^{-4}$, with the slight delta representing the selection penalty $\Delta$ derived in Appendix A.C.
 TABLE III: Diversity Gain Analysis ($\mathrm{SNR}=16$ dB)
@@
 4 (BC-TAS) $\mathbf{1.75\times 10^{-4}}$ $\mathbf{604.7\times}$
-As detailed in Table III, this massive improvement stems from the transition from a diversity order of $L=1$ (SISO) to an achieved diversity order of $L=N_{T}=4$.
-The table demonstrates that the reliability gain grows exponentially as additional spatial paths are exploited.
-In Rayleigh fading, the outage probability for selection diversity scales as $(P_{\mathrm{out,SISO}})^{N_{T}}$; thus, the observed three-order-of-magnitude reduction is consistent with theoretical expectations for a four-antenna selection system.
-By effectively ”ignoring” subcarriers in deep fades, BC-TAS ensures that the signal remains above the decoding threshold almost $99.99\%$ of the time, providing a robust backscatter-aware link without increasing hardware complexity.
-V-G Backscatter Crest Factor (BCF) and Link Stability
-While BER and outage probability characterize the primary link, the stability of the reflected field at the passive tag is governed by the BCF.
-As derived in Lemma 1, the BC-TAS framework utilizes a field-flattening cost term to minimize the temporal peaks of the backscatter signal.
-To ensure statistical rigor, we performed $1,000$ Monte Carlo realizations for each antenna configuration ($N_{t}\in\{2,4,8,16,32\}$).
-Furthermore, to mitigate the impact of imperfect CSI and prevent erratic antenna switching, we integrated a scalar Kalman Filter to smooth the estimated channel gains before selection.
+VI-G BCF and Link Stability
+The stability of the reflected field at the passive tag is governed by the BCF.
+As derived in Lemma 1, BC-TAS utilizes a field-flattening cost term to minimize temporal peaks.
+To ensure statistical rigor, we performed $1,000$ Monte Carlo trials per antenna configuration ($N_{t}\in\{2,\dots,32\}$), integrating a scalar Kalman Filter to smooth channel gains and prevent noise-driven antenna switching.
 Figure 4: Average BCF vs. $N_{t}$ across $1,000$ Monte Carlo trials in TGn Model B.
 As illustrated in Fig.
-4, the proposed BC-TAS (Kalman) model exhibits a superior downward trend in BCF compared to all baselines.
-At $N_{t}=32$, BC-TAS achieves a BCF of $\approx 14.0$ dB, representing a $1.1$ dB improvement over the OES and a $2.5$ dB improvement over the industry-standard NBAS.
-The proposed BC-TAS (Kalman) achieves a monotonic reduction in signal peaks, outperforming OES and SOTA baselines.
-This gap is significant because OES [SOTA_3_Exhaustive_Search], while optimal for the primary receiver’s BER, is ”tag-blind” and selects antennas that create constructive interference peaks at the tag’s location.
-In contrast, BC-TAS intelligently sacrifices marginal SNR gains to maintain a ”flatter” reflected field.
-The inclusion of the Kalman filter is vital here; by providing a smoothed estimate of the channel state, it prevents the high BCF values associated with rapid, noise-driven antenna switching across adjacent subcarriers.
-V-H PA Distortion Analysis
-To evaluate the impact of non-linear hardware impairments, we employ the Rapp PA model with a smoothness factor $p=2$ and a saturation voltage $V_{sat}=1$.
-The distortion is quantified through Error Vector Magnitude (EVM) and spectral compliance with the IEEE 802.11n mask.
+4, BC-TAS (Kalman) exhibits a superior downward trend in BCF compared to all baselines.
+At $N_{t}=32$, BC-TAS achieves a BCF of $\approx 14.0$ dB, a $1.1$ dB improvement over OES and $2.5$ dB over NBAS.
+While OES [SOTA_3_Exhaustive_Search] is optimal for BER, it remains “tag-blind,” often selecting antennas that create constructive interference peaks at the tag.
+Conversely, BC-TAS intelligently sacrifices marginal primary SNR to maintain a flatter reflected field.
+The Kalman filter is vital here; by smoothing channel state estimates, it suppresses the high BCF values typically caused by erratic, noise-driven antenna switching across subcarriers.
+The 1.1 dB BCF advantage over OES confirms Lemma 1, demonstrating that selection diversity can be dual-purposed for link reliability and peak power suppression.
+Furthermore, the 1,000 Monte Carlo realizations confirm that this is a persistent statistical property of the BC-TAS heuristic rather than a transient benefit.
+VI-H PA Distortion Analysis
+Hardware impairments are modeled using the Rapp PA ($p=2,V_{sat}=1$), with distortion quantified via EVM and IEEE 802.11n spectral mask compliance to ensure coexistence with legacy narrowband WSN nodes.
 Figure 5: Smoothed EVM performance vs. IBO highlighting the 16-QAM compliance threshold.
-As illustrated in the smoothed analysis in Fig.
-5, the proposed BC-TAS maintains a consistent linearity advantage over the MMSE MU-MIMO baseline across the entire operational range.
-While the MMSE baseline remains significantly above the 16-QAM EVM limit of $5.6\%$ throughout the saturation-to-linear transition, the proposed BC-TAS exhibits a steep descent toward compliance as the IBO increases.
-Specifically, BC-TAS approaches the required $5.6\%$ threshold at an IBO of approximately $13$ dB, demonstrating a significant reduction in constellation distortion compared to MU-MIMO.
-Furthermore, the indoor spectral compliance analysis in Fig.
-8 confirms that the BC-TAS architecture successfully suppresses out-of-band emissions.
-Even at an IBO of $10$ dB, the normalized PSD remains strictly within the 802.11n spectral mask limits, with the selection diversity effectively mitigating the spectral regrowth typically associated with high-PAPR OFDM signals.
-By operating the PA with less back-off while maintaining signal integrity, BC-TAS provides a double benefit: it maximizes the average transmit power available for the backscatter tag’s energy harvesting while simultaneously ensuring robust coexistence with legacy primary receivers in the 2.4/5 GHz bands.
-V-I Hardware Performance and Complexity Analysis
-The hardware feasibility of the proposed BC-TAS is evaluated by analyzing its robustness against PA non-linearities and its overall energy footprint compared to conventional MU-MIMO baselines.
-V-I1 PAPR Reduction and Statistical Analysis
-The PAPR is a critical bottleneck for PA efficiency in multi-antenna systems.
-We evaluate the statistical distribution of these peak power events using the Complementary Cumulative Distribution Function (CCDF), which defines the probability that the PAPR of an OFDM symbol exceeds a specific threshold $\gamma_{0}$.
+As shown in Fig.
+5, BC-TAS maintains a consistent linearity advantage over MMSE MU-MIMO.
+While the baseline remains above the 5.6% EVM limit for 16-QAM throughout most of the transition, BC-TAS achieves compliance at an IBO of approximately 13 dB.
+This reduction in constellation distortion is complemented by the spectral analysis in Fig.
+8.
+Even at 10 dB IBO, the normalized PSD remains strictly within IEEE 802.11n limits, as selection diversity mitigates the spectral regrowth typical of high-PAPR signals.
+By operating with less back-off while maintaining signal integrity, BC-TAS maximizes transmit power for tag energy harvesting while ensuring robust coexistence in the 2.4/5 GHz bands.
+VI-I Hardware Performance and Complexity Analysis
+The hardware feasibility of BC-TAS is evaluated by analyzing its robustness against PA non-linearities and its energy footprint compared to conventional MU-MIMO baselines.
+VI-I1 PAPR Reduction and Statistical Analysis
+PAPR remains a critical bottleneck for PA efficiency.
+We evaluate peak power events using the Complementary Cumulative Distribution Function (CCDF), defining the probability that an OFDM symbol’s PAPR exceeds a threshold $\gamma_{0}$.
 Figure 6: PAPR Reduction Analysis: Proposed BC-TAS vs. Conventional MMSE MU-MIMO highlighting the selection diversity gain.
-As illustrated in Fig.
-6, the proposed BC-TAS architecture exhibits a profound statistical advantage over conventional multiplexing-based methods.
-While both systems converge at low PAPR thresholds where average signal behavior dominates, they diverge significantly in the high-power tail region.
-At a clipping probability of $10^{-2}$ (1%), the proposed BC-TAS maintains a PAPR of approximately 7.1 dB, whereas the conventional MMSE MU-MIMO baseline reaches 9.5 dB.
-This significant 2.4 dB reduction in peak power is a direct consequence of the selection-based architecture.
-By leveraging the backscatter link to intelligently select subcarrier-antenna mappings that minimize constructive interference, BC-TAS effectively ”truncates” the probability of extreme peak events.
-This reduction in the PAPR tail is the primary driver for the improved hardware linearity and energy harvesting efficiency observed in subsequent analyses.
-V-I2 Energy Efficiency (EE)
-Energy Efficiency is defined as the ratio of spectral efficiency to the total power consumption, including RF chain activation and computational costs ($P_{RF}+P_{SEL}$).
+As shown in Fig.
+6, BC-TAS exhibits a profound statistical advantage.
+At a $10^{-2}$ clipping probability, BC-TAS maintains a PAPR of $\approx$7.1 dB, while the MMSE MU-MIMO baseline reaches 9.5 dB.
+This 2.4 dB reduction results from the selection-based architecture; by intelligently mapping subcarriers to antennas that minimize constructive interference, BC-TAS effectively “truncates” extreme peak events, driving the improved linearity and energy harvesting observed in subsequent analyses.
+VI-I2 Energy Efficiency (EE)
+EE is the ratio of spectral efficiency to total power consumption, including RF chain activation and computational costs ($P_{RF}+P_{SEL}$).
 Figure 7: System Energy Efficiency vs. PA Operating Point (Input Back-Off).
@@
 Computational Cost $\mathcal{O}(N_{t})$ $\mathcal{O}(N_{t}^{3})$ Significant
-Our results in Fig.
-7 demonstrate that BC-TAS provides a massive energy efficiency advantage, peaking at 0.075 bits/Hz/mW.
-In contrast, the MMSE baseline remains below 0.01 bits/Hz/mW due to the simultaneous activation of $N_{t}$ RF chains and high-complexity matrix inversions.
-V-I3 Spectral Mask Compliance
-Using the Rapp non-linearity model ($\rho=2$), we evaluate spectral regrowth against the IEEE 802.11n 20 MHz spectral mask.
-Figure 8: Indoor Spectral Compliance: BC-TAS Transmit Spectrum under IEEE 802.11n Mask.
-The normalized Power Spectral Density (PSD) in Fig.
-8 confirms that BC-TAS is fully compliant with the 802.11n mask even at low IBO levels.
-Sideband emissions are suppressed below 45-45 dB, allowing the transmitter to operate closer to the PA’s saturation region to maximize harvestable power for backscatter tags.
-V-J Joint Optimization and Fundamental Trade-off
-The flexibility of the BC-TAS framework is showcased through the fundamental trade-off between communication reliability (Primary Link BER) and the interference suppression benefit ($\Delta I$) afforded to the victim receiver.
+Results in Fig.
+7 show BC-TAS provides a massive EE advantage, peaking at 0.075 bits/Hz/mW.
+Conversely, the MMSE baseline remains below 0.01 bits/Hz/mW due to the simultaneous activation of $N_{t}$ RF chains and high-complexity matrix inversions.
+VI-I3 Spectral Mask Compliance
+Using the Rapp model ($p=2$), we evaluate spectral regrowth against the IEEE 802.11 mask.
+Figure 8: Indoor Spectral Compliance: BC-TAS Transmit Spectrum under IEEE 802.11 Mask.
+The normalized PSD in Fig.
+8 confirms BC-TAS compliance even at low IBO.
+Sideband emissions are suppressed below 45-45 dB, allowing the PA to operate nearer saturation to maximize harvestable power for backscatter tags while ensuring coexistence with legacy nodes.
+This creates a performance loop: selection diversity reduces PAPR, forcing the signal into the PA’s linear region and liberating power to increase tag harvesting efficiency $\eta_{H}$.
+By reducing the IBO while maintaining 16-QAM compliance, BC-TAS closes the gap between hardware constraints and passive link power requirements.
+VI-J Joint Optimization and Fundamental Trade-off
+The BC-TAS framework showcases a fundamental trade-off between primary link reliability (BER) and victim interference suppression ($\Delta I$).
 Figure 9: Fundamental Pareto trade-off between primary link reliability and victim interference suppression in TGn Model F.
 As illustrated in Fig.
-9, the system exhibits a clear operating boundary with a distinct ”knee” that defines the limits of selection diversity.
-The smoothed boundary (Savitzky-Golay filtered) reveals the optimal operating ”knee” where significant suppression is achieved with minimal BER penalty.
-By increasing the victim weight $\lambda_{V}$, the system achieves a maximum interference suppression of approximately $6.5$ dB.
-The ”Optimal Operating Point” in the figure is identified at $\Delta I\approx 4.5$ dB, where the system achieves a significant $2.8\times$ reduction in linear interference power while maintaining a primary link BER on the order of $10^{-3}$.
-This point represents the threshold of maximum marginal utility.
-In the region prior to this point ($0\leq\Delta I\leq 4.5$ dB), the BC-TAS algorithm successfully identifies antenna-subcarrier mappings that are naturally favorable to the victim with negligible impact on the primary link.
-However, beyond this $4.5$ dB threshold, the system enters a regime of diminishing returns; further gains in interference mitigation force the selection of antennas with significantly lower channel gains for the primary receiver, resulting in an exponential increase in BER as the diversity gain is exhausted to satisfy the strict coexistence constraint.
+9, the system exhibits a clear operating boundary with a distinct “knee” defining the limits of selection diversity.
+By increasing the victim weight $\lambda_{V}$, maximum suppression reaches $\approx 6.5$ dB.
+The optimal operating point is identified at $\Delta I\approx 4.5$ dB, where the system achieves a $2.8\times$ reduction in linear interference power while maintaining a primary link BER of $10^{-3}$.
+In the region $\Delta I\leq 4.5$ dB, BC-TAS identifies antenna-subcarrier mappings naturally favorable to the victim with negligible BER penalty.
+Beyond this threshold, the system enters a regime of diminishing returns where further interference mitigation forces the selection of significantly weaker primary channel gains, resulting in an exponential BER increase.
+The empirical 6.5 dB maximum suppression aligns with the $10\log_{10}(N_{t})\approx 6.02$ dB theoretical bound, validating the “PDF squeezing” effect.
+This confirms that the SDR expansion is a direct consequence of lowering the interference floor $\mathbb{E}[P_{B}|\lambda_{T}]$, allowing the SAW sensor to operate in a cleaner spectral environment.
 Figure 10: Pareto front scaling with antenna count $N_{tx}\in\{2,4,8\}$.
-The impact of hardware complexity on this trade-off is further explored in Fig.
+The impact of hardware complexity is explored in Fig.
 10.
-As $N_{tx}$ increases, the Pareto front shifts the entire operating boundary toward the high-suppression, low-BER region..
-This shift indicates that higher-order selection diversity provides a more favorable trade-off surface, allowing for simultaneous improvements in both reliability and coexistence.
-For $N_{tx}=8$, the system can achieve suppression levels exceeding $7$ dB while remaining well below the BER levels of a $N_{tx}=2$ system with zero suppression.
-V-K Spatial Correlation, PAPR, and Backscatter Efficiency
-Conventional multi-user MIMO (MU-MIMO) systems and Space-Time Block Coding (STBC) schemes face significant challenges in WSN deployments, particularly regarding PAPR and backscatter energy harvesting efficiency.
-While linear receivers like MMSE detectors experience severe noise enhancement under high transmit-side spatial correlation ($\mathbf{R}_{\mathrm{tx}}$), STBC requires multiple active RF chains, which increases hardware cost and limits the power available for backscatter modulation.
-In contrast, the proposed BC-TAS scheme addresses these bottlenecks by exploiting selection diversity across spatial and frequency domains.
-By activating only a single transmit antenna per subcarrier, BC-TAS inherently solves the high PAPR issue typically found in multi-stream MIMO, as the localized power concentration reduces the dynamic range of the transmitted signal.
-Furthermore, by identifying the most favorable propagation paths that maximize backscatter reflection power while minimizing interference, the scheme ensures high energy efficiency for passive sensor nodes.
+As $N_{tx}$ increases, the Pareto front shifts toward the high-suppression, low-BER region, indicating that higher-order selection diversity provides a more favorable trade-off surface.
+For $N_{tx}=8$, the system achieves suppression exceeding $7$ dB while remaining significantly more reliable than an $N_{tx}=2$ system with zero suppression.
+VI-K Spatial Correlation, PAPR, and Backscatter Efficiency
+Conventional MU-MIMO and Space-Time Block Coding (STBC) schemes face challenges in WSNs regarding PAPR and energy harvesting efficiency.
+While MMSE detectors suffer noise enhancement under high spatial correlation ($\mathbf{R}_{\mathrm{tx}}$) and STBC requires multiple costly RF chains, BC-TAS exploits selection diversity to bypass these bottlenecks.
+By activating a single antenna per subcarrier, BC-TAS reduces signal dynamic range, inherently solving the multi-stream PAPR issue while maximizing backscatter reflection power.
 Figure 11: Reliability robustness vs. transmit antenna correlation $\mathbf{R}_{\mathrm{tx}}$ at $\mathrm{SNR}=16$ dB.
-As illustrated in Fig.
-11, at a fixed SNR of $16$ dB, the MMSE baseline exhibits an exponential rise in BER, reaching approximately $2.8\times 10^{-2}$ at $\mathbf{R}_{\mathrm{tx}}=0.9$.
-While STBC provides better stability through orthogonal diversity, it maintains a performance gap significantly wider than the proposed scheme.
-The proposed BC-TAS demonstrates superior robustness, achieving a BER of approximately $4.5\times 10^{-4}$ at extreme correlation levels ($0.9$).
-This result confirms that BC-TAS effectively circumvents the ill-conditioned channel matrices that plague MMSE and the hardware overhead of STBC.
-By focusing transmit energy into a single optimal antenna-subcarrier mapping, BC-TAS maximizes the signal-to-noise ratio for both the primary receiver and the backscatter tag, providing a low-complexity solution for high-density, correlated WSN environments.
-V-L Physical Layer Insights: Frequency-Selective Notching
-To provide physical intuition into the BC-TAS mechanism, we analyze the Spectral Power Density (SPD) of the transmitted signal as observed at the victim receiver in a frequency-selective TGn Model F environment.
+As shown in Fig.
+11, at 16 dB SNR, the MMSE baseline exhibits an exponential BER rise, reaching $2.8\times 10^{-2}$ at $\mathbf{R}_{\mathrm{tx}}=0.9$.
+Although STBC offers better stability, BC-TAS demonstrates superior robustness, achieving a BER of $\approx 4.5\times 10^{-4}$ at the same extreme correlation.
+This confirms that BC-TAS circumvents the ill-conditioned matrices inherent in MMSE and the hardware overhead of STBC.
+By focusing transmit energy into optimal antenna-subcarrier mappings, BC-TAS maximizes the SNR for both the primary receiver and the backscatter tag, providing a low-complexity solution for high-density, correlated WSN environments.
+VI-L Physical Layer Insights: Frequency-Selective Notching
+To provide physical intuition, we analyze the Spectral Power Density (SPD) at the victim receiver in a frequency-selective TGn Model F environment.
 Figure 12: SPD Analysis: Frequency-selective notching at the victim receiver.
 As illustrated in Fig.
-12, when the system operates in ”Protective Mode” ($\lambda_{V}=10$), the algorithm selectively maps subcarriers to antennas that exhibit deep fading at the victim’s spatial location.
-This results in a ”notching” effect where the interference power is surgically removed from specific subcarriers.
-For $\lambda_{V}=10$, the BC-TAS algorithm identifies and exploits natural channel nulls across the subcarrier index to protect the victim.
-Crucially, this observed ”notching” validates the spatial-frequency decoupling inherent in our algorithm.
-It proves that the interference nulling is achieved through channel-aware selection diversity rather than expensive digital filtering or active cancellation hardware—computational overheads that a passive backscatter tag cannot support.
-By exploiting the natural multipath environment to ”hide” the signal from the victim, the BC-TAS framework justifies its low hardware complexity while maintaining high primary link power on subcarriers where the victim is naturally shielded.
-VI Discussion and Theoretical Validation
-In this section, we synthesize our numerical findings with the analytical framework derived in Section III and Appendix A.
-The alignment between the observed gains and the order-statistics model provides a robust validation of the BC-TAS framework.
-VI-A Linking Field Flattening to Lemma 1
-The results in Fig.
-4 provide empirical confirmation of the ”Field Flattening” effect described in Lemma 1.
-The $1.1$ dB BCF advantage over OES at $N_{t}=32$ demonstrates that selection diversity can be dual-purposed for both link reliability and peak power suppression.
-The use of 1,000 Monte Carlo trials confirms that this is not a transient benefit but a persistent statistical property of the BC-TAS heuristic.
-By smoothing the measurement noise through the Kalman update (Section III-C), the system effectively ”regularizes” the cost function, ensuring that the selected antenna-subcarrier mapping represents a stable propagation path rather than a noisy outlier.
-This stability is critical for low-complexity backscatter tags that lack the processing power to track high-variance reflected waveforms.
-VI-B Linking Interference Suppression to SDR Expansion
-The observed $2.4$ dB PAPR reduction and the corresponding interference suppression ”knee” in Fig.
-9 directly validate the result in Lemma 1.
-As established analytically, the selection of the minimum interference path among $N_{t}=4$ antennas yields an expected power reduction scaling by $1/N_{t}$.
-In the logarithmic domain, this theoretical shift is approximately $10\log_{10}(4)\approx 6.02$ dB.
-Our empirical maximum suppression of $6.5$ dB (see Fig.
-9) aligns closely with this bound.
-The expansion of the SDR is therefore a direct consequence of this ”PDF squeezing” effect, where the background interference floor $\mathbb{E}[P_{B}|\lambda_{T}]$ is lowered, allowing the SAW sensor to operate in a significantly cleaner spectral environment.
-VI-C Validation of the three-order-of-magnitude Outage Reduction
-The massive reliability improvement reported in Table III is mathematically consistent with the diversity order $L=N_{t}$.
-Theoretically, the outage probability for selection diversity follows $P_{out}\approx(P_{out,SISO})^{N_{t}}$.
-Given our SISO baseline outage of $0.105$ (at $16$ dB), the theoretical prediction for $N_{t}=4$ is $(0.105)^{4}\approx 1.2\times 10^{-4}$.
-Our simulated BC-TAS outage probability of $1.75\times 10^{-4}$ tracks this theoretical floor with high precision.
-The slight delta is attributed to the ”Selection Penalty” $\Delta$, where the system occasionally sacrifices the absolute maximum SNR to satisfy the weight constraints $\lambda_{T}$ and $\lambda_{V}$, as derived in Appendix A.C.
-VI-D The PAPR-Linearity-Efficiency Loop
-The $2.4$ dB gain in the PAPR CCDF (Fig.
-6) serves as the hardware-level justification for the EVM and Energy Efficiency gains.
-By selecting antenna mappings that truncate the high-power tail of the OFDM distribution, BC-TAS forces the signal into the linear region of the Rapp PA model.
-This enables the transmitter to reduce the IBO by approximately $3$ dB while maintaining 16-QAM compliance (see Fig.
-5).
-This ”liberated” power is redirected toward tag illumination, increasing the harvesting efficiency $\eta_{H}$.
-Consequently, the system closes a performance loop: selection diversity reduces PAPR, which improves PA efficiency, which ultimately expands the power budget for the passive backscatter link.
+12, operating in ”Protective Mode” ($\lambda_{V}=10$) forces the algorithm to map subcarriers to antennas exhibiting deep fading at the victim’s location.
+This creates a ”notching” effect where interference power is surgically removed from specific subcarriers by exploiting natural channel nulls.
+Crucially, this validates the spatial-frequency decoupling inherent in BC-TAS: interference nulling is achieved through channel-aware selection rather than expensive digital filtering or active cancellation hardware.
+By utilizing the multipath environment to ”hide” the signal from the victim, BC-TAS maintains low hardware complexity while concentrating transmit power on subcarriers where the victim is naturally shielded.
 VII Conclusion
@@
 In a conventional Max-SNR TAS system, the selected gain $G_{\max}=\max\{G_{1},\dots,G_{N_{t}}\}$ follows the harmonic expectation:
-$\mathbb{E}[G_{\max}]=\sigma_{L}^{2}\sum_{n=1}^{N_{t}}\frac{1}{n}\approx\sigma_{L}^{2}\ln(N_{t})+\gamma_{e}$ (13)
+$BC-TAS\mathbb{E}[G_{\max}]=\sigma_{L}^{2}\sum_{n=1}^{N_{t}}\frac{1}{n}\approx\sigma_{L}^{2}\ln(N_{t})+\gamma_{e}$ (13)
 where $\gamma_{e}$ is the Euler-Mascheroni constant. This serves as the theoretical upper bound for primary link performance.